### v1

# supSupplementary References

Abstract
Systems that answer questions by reviewing the scientific literature are becoming increasingly feasible. To draw reliable conclusions, these systems should take into account the quality of available evidence, placing more weight on studies that use a valid methodology. We present a benchmark for measuring the methodological strength of biomedical papers, drawing on the risk-of-bias framework used for systematic reviews. The four benchmark tasks, drawn from more than 500 papers, cover the analysis of research study methodology, followed by evaluation of risk of bias in these studies. The benchmark contains 2000 expert-generated bias annotations, and a human-validated pipeline for fine-grained alignment with research paper content. We evaluate a range of large language models on the benchmark, and find that these models fall significantly short of expert-level performance. By providing a standardized tool for measuring judgments of study quality, the benchmark can help to guide systems that perform large-scale aggregation of scientific data. The dataset is available at https://github.com/RoBBR-Benchmark/RoBBR.
1 Introduction
Systems that automatically answer questions by reviewing the scientific literature are becoming increasingly feasible. These systems offer the potential to provide scientists with on-demand access to knowledge that is synthesized from across the literature. This can help scientists understand what is already known about topics outside of their research focus, and help clinicians and practitioners keep up to date with best practices.
When assessing what is known about a field, not all studies should be weighed equally I et al. [2023]. Studies with stronger methodologies contribute more to a body of evidence than those with weaker methodologies. By weighing studies appropriately, systems can increase the reliability of their summaries and recommendations Turner et al. [2009].
In biomedical research, there is a large body of work which investigates the factors that decrease the validity of a study’s methods, and best practices for addressing these issues I et al. [2023], Welton et al. [2009]. Reporting bias occurs when there are systematic differences in how outcomes are reported or disclosed between the groups that are compared (e.g., the results of the treatment group are more frequently or favorably published than those in the placebo group). This bias is often best addressed by adhering to strict reporting standards and protocols, such as registering the study in advance and committing to publish all results Kotz and West [2022]. Other biases, such as detection bias, attrition bias, and selection biases, can frequently be addressed using careful study design and implementation Berger [2005], JPT et al. [2023], I et al. [2023].
This paper presents the RoBBR benchmark for assessing the methodological strength of biomedical studies. The benchmark follows the risk of bias framework developed by Cochrane Jpt et al. [2023], JPT et al. [2023], Higgins et al. [2011], Practice and of Care [EPOC], Sterne et al. [2014], a global independent network that conducts systematic reviews of healthcare interventions. Cochrane’s risk of bias framework is widely recognized as a standard for evaluating the quality and reliability of research studies. It serves as an important resource for national health agencies in various countries, including the United States Viswanathan et al. [2012], the United Kingdom Alderson and Tan [2011], and Australia NHMRC [2019]. Notably, Cochrane’s systematic reviews are used to formulate national and international guidelines on healthcare practices Bunn et al. [2015]. Furthermore, empirical comparisons between Cochrane and non-Cochrane systematic reviews consistently demonstrate higher methodological quality in those conducted by Cochrane, highlighting its critical role in advancing evidence-based healthcare Goldkuhle et al. [2018], Page et al. [2018], Matthias et al. [2020], Tsoi et al. [2020].
Figure 1: The figure illustrates the components of a risk of bias assessment. Given a bias category and a research study, a meta-reviewer identifies factors that make the study high or low risk for that bias. The meta-reviewer summarizes their findings in a support judgment, and then make a final decision about the risk level.
2 Background
The RoBBR benchmark aims to provide an expert-validated tool for assessing whether models can make risk of bias judgments for biomedical papers. The benchmark uses a risk-of-bias framework developed in the context of systematic reviews, which is the gold standard for assessing biomedical studies. We provide more background on systematic reviews and the risk of bias framework below.
Systematic reviews and meta-analyses are tools used in evidence-based medicine, which synthesize evidence from multiple studies to answer important clinical questions. They help researchers and healthcare professionals make informed decisions based on the best available evidence Ahn and Kang [2018], TJ et al. [2023], JJ et al. [2023].
A systematic review is a comprehensive search of the literature to identify all relevant studies on a particular topic TJ et al. [2023], C et al. [2023]. The review process follows a pre-specified protocol to minimize bias and ensure transparency. The protocol outlines the research question, search strategy, study selection criteria, data extraction methods, and plans for quality assessment and synthesis of result Jpt et al. [2023], Page et al. [2021].
A meta-analysis is a type of systematic review that uses statistical methods to combine the results from multiple studies Ahn and Kang [2018]. By pooling the data across studies, a meta-analysis can provide a more precise estimate of the treatment effect JJ et al. [2023], Guevara et al. [2004] and explore potential sources of heterogeneity Higgins and Thompson [2002].
The research question for a systematic review is often structured using the PICO framework, which stands for Participant, Intervention, Comparison, and Outcome Aslam and Emmanuel [2010]. The PICO elements help define the key components of the research question and guide the study selection process J et al. [2023], JE et al. [2023a]. For example, a systematic review on the effectiveness of a new hypertension drug might have the following PICO: Participant: Adults with primary hypertension; Intervention: New antihypertensive drug; Comparison: Placebo or standard care; Outcome: Change in systolic and diastolic blood pressure.
By clearly defining the PICO elements, the review authors can develop a focused search strategy and establish explicit criteria for selecting studies to include in the review J et al. [2023], JE et al. [2023b].
Inclusion criteria typically specify the characteristics of the study population, the interventions and comparisons of interest, the outcomes to be measured, and the study designs to be included JE et al. [2023b], Page et al. [2021], Tawfik et al. [2019] (e.g., randomized controlled trials). Exclusion criteria describe reasons why a study must be excluded from an analysis, such as using the wrong study design C et al. [2023], Page et al. [2021], Tawfik et al. [2019].
Risk of bias assessment involves evaluating each included study to determine the likelihood that its results may be biased JPT et al. [2023], Practice and of Care [EPOC], Sterne et al. [2014]. The Cochrane tool for assessing risk of bias Higgins et al. [2011] is widely used and assesses several types of bias, including random sequence generation, allocation concealment, blinding of participants and outcome assessors, incomplete outcome data, selective reporting, and other sources of bias. During the assessment, two review authors independently evaluate the risk of bias using the Cochrane tool and resolve disagreements through discussion. They record their rationale for each bias assessment. These rationales are referred to as "support judgments."
For example, a support judgment of Reporting Bias could be "The study protocol was registered (ClinicalTrials.gov NCT01863706) but not all of the outcomes projected by methodological descriptions were reported as results in the study report (cases of diarrhoea, nausea and vomiting were not completely reported). Moreover, the study publication reports outcomes (hypotension, nausea, transfusion) not listed in the registered protocol." A decision regarding the risk of reporting bias is then made on the basis of the support judgment. See Figure 1 for an illustration.
Studies with a high risk of bias may overestimate or underestimate the true treatment effect, leading to inaccurate conclusions in the systematic review. In this paper, we consider eight categories of biases: selection bias, attrition bias, performance bias, detection bias, reporting bias, analysis bias, special bias (specific to individual meta-analyses), and other bias (i.e. biases not covered by the other categories). We benchmark model performance on assessing papers for each of these bias categories.
3 Benchmark Development
The RoBBR benchmark is derived from 63 Cochrane meta-analyses and 532 papers referenced in those meta-analyses. All of the meta-analyses and papers have a Creative Commons license or are public domain. The benchmark is designed to evaluate a model’s ability to judge the risk of bias in biomedical reports.
The data were collected in April 2024. We included all meta-analyses from Cochrane that are in the BioC database, are classified as intervention reviews, have a CC-BY-NC license, and contribute at least one data point for each task. We obtained the XML format of the papers included in these meta-analyses either through the BioC API or by manually downloading and converting PDFs to XML using GROBID. We then extract task data from the XML of these papers.
The benchmark is divided into four main tasks (See in Appendix A.3 for detailed statistics):
  • Study Inclusion/Exclusion (Task 1): Determine whether a study’s objectives and methodology fit a set of requirements.
  • Bias Retrieval (Task 2): Retrieve sentences from the paper that support a judgment about the study’s risk of bias.
  • Support Judgment Selection (Task 3): From a list of options, select the judgment that best supports a determination of the study’s risk of bias.
  • Risk Level Determination (Task 4): Assess the study’s risk of specific biases.
These four tasks mirror the procedure followed by meta-reviewers. Initially, reviewers decide if a study should be included or excluded based on predefined criteria, which reflect the goals of the review (Task 1). Included studies will be analyzed for indications of the risk of bias (Task 2). Reviewers will then summarize the evidence collected from the study, and analyze its implications (Task 3). Finally, they assess the risk level of specific biases for each identified bias (Task 4).
By structuring the dataset around these tasks, we provide a comprehensive framework for simulating the nuanced process of risk of bias assessment in meta-analysis.
3.1 Train/Test Split
We randomly select 23 meta-analyses and their 228 corresponding papers in the dev set. We put the other 40 meta-analyses and their corresponding 304 papers in the test set.
Figure 2: Study Inclusion/Exclusion. In Task 1, the model sees a search protocol and objective, and is asked to determine whether a study’s methodology fits those criteria.
3.2 Task 1: Study Inclusion/Exclusion
For each potentially relevant study to be included in a meta-analysis, the search protocol information and meta-analysis objective are provided as input, and the task is to decide if the study should be included in the meta-analysis. The objective is included in order to provide similar context to what is normally available for a meta-reviewer and make it easier to interpret the criteria in the search protocol. An example is provided in Figure 2.
3.3 Task 2: Risk of Bias Sentence Retrieval
When judging the risk of bias for a specific study, the review authors provide support judgments which justify their risk of bias rating. These support judgments are grounded in specific aspects of the study, which the support judgment describes. Task 2 aims to test a model’s ability to retrieve the correct source for the support judgment from a paper’s text. The support judgments are intentionally omitted in the Bias Retrieval task. This exclusion simulates the systematic review process by challenging the model to independently generate these justifications. Including them would come close to providing the answer, undermining the purpose of the task. See Figure 3 for an illustration.
Some review authors directly quote sentences from the paper in their support judgment, allowing for straightforward extraction and matching of these sentences with the paper’s text. However, most reviewers prefer to paraphrase the sentences or engage in deeper analysis, synthesizing information from multiple parts of the paper.
To match these judgments with sentences from the paper, we use a three part annotation pipeline.
First, a support judgment often contains multiple pieces of information. We decompose the support judgment into distinct, non-overlapping pieces, each focusing on a specific aspect of the original support judgment.
Second, we exclude any part of the support judgment containing a negation or non-specific commentary, because these parts of the judgment cannot be matched to any sentence. The remaining parts of the support judgments generally contain specific claims about the study.
Third, these aspects are mapped to specific sentences within the paper, allowing us to pinpoint the exact sources of information that underlie each aspect of the judgment. A single aspect could potentially be supported by multiple sentences from the paper, and similarly a single sentence can support multiple aspects.
Each of these steps, including the aspect decomposition, uses GPT-4 (gpt-4-0125) OpenAI [2023]. The prompts can be found in the Appendix C.
We manually inspect the 2221 support judgment aspects generated in the second stage, to ensure that all aspects are related to specific information from the paper.
In the next section, we evaluate annotation quality by comparing to human judgments.
Figure 3: Risk of Bias: Sentence Retrieval. In Task 2, the goal is to retrieve sentences from the paper (the Candidate Pool) which support a risk of bias judgment. Performance is evaluated against a Cochrane review support judgment, which has been split into aspects.
3.3.1 Evaluating LLM Annotations
To match each aspect from a support judgment with specific sentences in the paper, we need to evaluate each sentence (on average, 200 sentences) against each aspect. For a support judgment with 5 aspects, this results in approximately 1,000 decisions. Given the scale of our entire dataset, which includes over 500 (bias, study) pairs, this translates to more than 1 million decisions – an impractical task for human annotators to perform manually.
We hypothesize that determining whether a sentence matches an aspect is straightforward enough for GPT-4 to achieve human-level performance with a specialized prompt. Following the protocol introduced by Wang et al. [2023], we created a development set of 30 papers. A single support judgment aspect was chosen for each paper, and every sentence in the paper was annotated for that aspect, totaling 8789 decisions. We optimized the instruction prompt for GPT-4 on this development set. Details about the prompt and experimental procedure are provided in the Appendix C.
Once the prompt was finalized, four annotators were tasked with annotating 50 randomly sampled (aspect, full paper) pairs (all annotators saw the same pairs). Every sentence in each paper was annotated, requiring 13575 annotations in total. The annotators were divided into two teams, with each team consisting of two annotators. Each annotator performed the annotation tasks individually given a pre-defined annotation guideline. Annotators from the same team then collaborated to resolve differences and eliminate mistakes. Refer to the Appendix B for the full annotation guidelines.
Each team produced a set of annotation results. GPT-4 annotated the same (aspect, paper) pairs. We calculated Binary Accuracy, F1-binary, Spearman Correlation, and Kappa Coefficient between the two human teams and between each human team and GPT-4. Table 1 shows that the human teams matched each other and GPT-4 more than 99% of the time. The other agreement metrics have values in the low-to-mid 70’s, indicating substantial agreement among annotators. [^1] These metrics are lower than exact agreement because of class imbalance. Positive matches between sentences and aspects are rare. Bootstrapped hypothesis tests do not demonstrate any statistical difference between GPT-4 and the human annotators.
Table 1: Inter-annotator agreement among the two teams of humans and the human teams and GPT-4.
Metrics Human & Human Average Human & GPT Average p-value
Exact Accuracy 99.4 $\pm$ 0.2 99.3 $\pm$ 0.2 0.31
F1 Binary 73.4 $\pm$ 6.3 71.7 $\pm$ 5.6 0.55
Cohen’s $\kappa$ 73.1 $\pm$ 6.3 71.4 $\pm$ 5.6 0.54
Spearman’s $\rho$ 73.8 $\pm$ 6.0 71.7 $\pm$ 5.6 0.43
3.4 Task 3: Risk of Bias Support Judgment Selection
The previous task evaluates a model’s ability to extract information from biomedical papers to in order to justify assessments of bias. However, this approach alone is insufficient for a risk-of-bias assessment. The evaluation must also include the model’s ability to reason about the retrieved information. For example, a paper might report the loss to follow-up yet fail to provide reasons for missing data, which raises the possibility of attrition bias. Similarly, while a paper may state that it implemented double-blinding, expert reviewers could consider such blinding infeasible due to the nature of the trial, potentially introducing performance bias.
To address this, we propose a multiple-choice task where the model selects the correct support judgment. This task includes three synthetically generated hard options, three options derived from other papers’ support judgments for the same bias category, and the Cochrane reviewer’s actual judgment. We prompt GPT-4 to generate the three synthetic options by imitating support judgments from other papers concerning the same type of bias. These options are tailored to be paper-specific while maintaining the underlying reasoning. See Appendix D for details on construction of these synthetic judgments.
In order to ensure that all incorrect options were actually incorrect, we performed filtering in two stages. GPT-4 was used to perform preliminary filtering (the prompt is shown in Figure 13 of Appendix D). In the second stage, we manually reviewed all of the remaining options, and removed any that were not incorrect.
Figure 4 shows an example of a multiple-choice question with one correct answer and three synthetically generated answers. Both options C and D refer to the same information from the paper, in this case the operational challenges affecting the trial’s methodology, but describe different reasoning given this information. This demonstrates that retrieving the correct information from the paper is not sufficient for solving this task.
Figure 4: Risk of Bias Support Judgment Selection. In Task 3, the model is shown a paper, and asked which of 7 support judgments provides the best explanation of the paper’s risk of bias. (Only four options are shown here for illustration purposes.)
3.5 Task 4: Risk Level Determination
The final task directly evaluates a model’s ability to assess the risk of bias in biomedical research papers. It requires the model to categorize the risk level of a specific bias as high, low, or unclear. The input to the model includes the whole paper, the PICO of the study, the objective of the meta-analysis, and the definition of the bias, as illustrated in Figure 5.
The review authors use bias with definitions from Cochrane Handbook 5.1 Higgins and Green [2011], Cochrane Effective Practice and Organisation of Care Practice and of Care [EPOC], A Cochrane Risk of Bias Assessment Tool Sterne et al. [2014], or specific biases that are defined within the meta-analysis. For the last category, we extract the bias definition from the specific meta-analysis. (Bias definitions are also extracted for Tasks 2 and 3.)
This assessment requires synthesizing various elements identified in the previous tasks, including the accurate retrieval and interpretation of relevant sentences and the ability to critically evaluate the sufficiency and reliability of the evidence presented regarding each bias.
Figure 5: Risk Level Determination. In Task 4, the goal is to provide an assessment of the paper’s risk of specific biases.
4 Experiments
4.1 Experimental Models and Procedures
We evaluate GPT-4o OpenAI [2024], Claude3-Opus Anthropic [2024a], Claude3.5-Sonnet Anthropic [2024b], and Gemini-1.5-Pro Google [2024]. Due to the long-context nature of RoBBR’s tasks, with many datapoints exceeding 16k tokens, many open-sourced LLMs AI [2024], Poli et al. [2023], Li et al. [2023] are considered but none of which is equipped to evaluate RoBBR.
We use chain-of-thought as prompting strategy to stabilize the model performance. We optimize our prompts on a separate development set. See Appendix E for details on prompt optimization.
4.2 Evaluation metrics
For Task 1, study inclusion/exclusion, we report macro-F1, accuracy, and weighted accuracy.
For Task 2, Risk of Bias Sentence Retrieval, we report aspect-level recall metrics, which measure the proportion of the bias aspects that are covered by the retrieved sentences. Specifically, Bias Recall (BR) @ 3 measures the proportion aspects covered when three sentences are retrieved. BR @ Optimal measures recall when the “optimal" number of sentences are retrieved. The optimal number of sentences is defined as the fewest which are sufficient to cover all bias aspects.
For Task 3, Risk of Bias Support Judgment Selection, we report accuracy.
For Task 4, Risk Level Determination, we report accuracy and weighted accuracy.
4.3 Categorization
For Tasks 2, 3, and 4, we consider 8 categories of biases: selection bias, attrition bias, performance bias, detection bias, reporting bias, analysis bias, special bias (the biases that are meta-analysis specific), and other bias (i.e. biases that do not fall into any of the other categories). In order to reduce evaluation costs, we randomly sampled 50 task instances for some categories. It is important to note that systematic reviews vary in which biases they report. When the authors judge that a bias is not important for the objective of a review, they may decide not to evaluate a specific bias. As a result, not all categories have the same number of examples, and some categories have fewer than 50 examples. See Table 10 for detailed statistics.
5 Results
Table 2: Results for Task 1: Study Inclusion/Exclusion
F1 Accuracy Weighted Accuracy
Claude3-Opus 75.5 77.1 74.9
Claude3-Sonnet 80.6 81.1 80.8
Gemini-1.5-Pro 72.2 72.7 73.0
GPT-4o 75.5 75.6 77.6
[…]
Bias Type
Model Full Selection Attrition Performance Detection Reporting Analysis Special Other
n = 337 n = 117 n = 57 n= 50 n = 50 n = 35 n = 13 n = 7 n = 23
Bias Recall@Optimal
Claude3-Opus 39.0 55.4 22.0 57.3 46.4 11.0 47.9 13.9 4.4
Gemini-1.5-Pro 41.1 57.2 28.7 49.1 48.8 23.3 54.2 13.9 3.3
GPT-4o 44.8 60.3 28.6 64.6 48.7 32.9 28.7 28.0 2.3
Bias Recall@3
Claude3-Opus 52.3 67.9 40.2 63.1 55.9 25.5 55.9 64.6 19.7
Gemini-1.5-Pro 56.1 69.7 44.1 73.7 65.3 35.6 56.3 35.5 6.7
GPT-4o 61.8 72.2 47.9 82.0 78.5 43.2 67.9 50.4 10.3
5.1 Performance on the Four Tasks
In Task 1 (Study Inclusion/Exclusion), shown in Table 2, Claude3-Opus and GPT-4o perform best in terms of F1 score. Claude3-Opus has higher accuracy and GPT-4o higher weighted accuracy.
In Task 2 (Sentence Retrieval Evaluation Results) 3, GPT-4o has stronger results than both Gemini-1.5 and Claude3-Opus for almost all bias types. However, in Task 3 (Support Judgment Selection Evaluation Results) 4, Claude3-Opus has the strongest results.
Tasks 2 and 3 are critical for evaluating the different capabilities of the LLMs. Task 2 is a long-context task, requiring retrieval of a non-redundant set of sentences from a 200-sentence paper. Task 3 requires complex reasoning to evaluate the retrieved sentences and identify the implications for risk of bias. Our results demonstrate that GPT-4o performs better in long-context retrieval tasks, while Claude3-Opus exhibits stronger reasoning abilities.
Figure 4 illustrates Claude3-Opus’s stronger reasoning skills. Options C and D present different reasoning based on the same information from the paper. Notably, GPT-4o is misled by a deceptive option, whereas Claude3-Opus correctly handles the multiple-choice question.
Task 4 (Risk Level Determination Evaluation Results) 5 provides an overall assessment of each model’s ability to determine the risk level, integrating all information from the previous tasks. The three models show comparable performance in this task.
[…]
Bias Type
Model Full Selection Attrition Performance Detection Reporting Analysis Special
n = 310 n = 86 n = 60 n = 52 n = 56 n = 50 n = 8 n = 6
Accuracy
Claude3-Opus 51.2 62.6 56.7 48.0 46.5 44.2 50.0 0.0
Claude3.5-Sonnet 60.2 69.8 79.9 55.8 44.8 51.8 62.5 0.0
Gemini-1.5-Pro 49.3 57.1 59.9 46.4 42.8 40.2 49.3 16.6
GPT-4o 45.8 52.3 63.1 42.5 37.7 28.0 74.5 0.0
[…]
Bias Type
Model Full Selection Attrition Performance Detection Reporting Analysis Special Other
n = 382 n = 93 n = 50 n = 50 n = 50 n = 50 n = 18 n = 26 n = 50
F1
Claude3-Opus 47.7 54.3 34.1 46.7 37.4 23.0 69.9 27.9 27.9
Claude3.5-Sonnet 49.7 62.7 23.9 49.9 46.7 30.1 48.3 28.7 33.8
Gemini-1.5-Pro 43.1 53.9 26.0 43.7 46.8 30.4 44.6 20.1 22.8
GPT-4o 51.9 59.0 36.5 48.9 36.0 27.7 53.8 34.7 29.8
Accuracy
Claude3-Opus 56.8 64.6 58.1 48.2 41.7 40.0 88.9 53.8 72.2
Claude3.5-Sonnet 62.3 64.4 52.0 52.1 55.9 64.3 83.0 69.5 72.1
Gemini-1.5-Pro 49.3 62.5 42.4 46.5 48.1 56.0 77.5 42.4 23.7
GPT-4o 62.6 71.0 68.3 54.4 40.1 57.6 88.7 65.4 62.1
Weighted Accuracy
Claude3-Opus 49.5 55.3 36.2 48.0 46.5 30.4 69.3 28.2 32.4
Claude3.5-Sonnet 50.3 63.1 24.1 51.6 48.2 32.1 46.8 36.1 36.1
Gemini-1.5-Pro 47.9 56.3 34.0 44.4 54.2 36.4 43.6 22.2 32.5
GPT-4o 53.6 60.4 36.3 51.3 40.7 30.0 50.0 64.3 32.6
5.2 Performance across Bias Types
The three models exhibit significant variations in performance across different biases due to the varying challenges associated with each bias type. All models have stronger performance at assessing selection bias, as biomedical papers typically detail their randomization processes explicitly.
In contrast, assessing reporting bias requires a comprehensive understanding of the entire paper to determine whether all outcomes specified in the study protocol are reported. Evaluating attrition bias involves extracting data on participant dropout during the trial and assessing how this dropout impacts the study’s statistical estimates. The most challenging bias type to assess is "other bias," where the model must grasp all potential biases, such as those stemming from funding sources or bias due to lack of power. Successfully identifying them requires a thorough understanding of all possible biases a study might have, thereby necessitating the treatment of each bias type as a distinct task.
6 Related Work
Scientific and Biomedical Information Retrieval has been the subject of a significant body of work in recent years Jin et al. [2023], MacAvaney et al. [2020], Wang et al. [2023]. A subset of this work has been focused on benchmarks Ben Abacha and Demner-Fushman [2019], Krithara et al. [2023], Voorhees et al. [2021], Cohen et al. [2017].
Study Screening Many works have focused on the task of screening studies for inclusion in systematic reviews Cohen et al. [2006], Khabsa et al. [2016], Kontonatsios et al. [2017], El-Gayar et al. [2015], with more recent work often focusing on LLMs Wilkins [2023], Tran et al. , Guo et al. [2024], Na et al. [2024], Cai et al. [2023], Ye et al. [2024], Wang et al. [2024]. Huotala et al. [2024] conduct an evaluation of LLM paper screening, finding that few-shot chain-of-thought prediction performs comparably to human annotators. Robinson et al. [2023] leverage an instruction-tuned Guanaco-7B model on a large dataset of systematic reviews to exceed the screening performance of ChatGPT. Cohen et al. [2006] produce a dataset of screening judgments for 15 systematic drug reviews.
Study Risk of Bias Risk of bias identification requires a thorough reading of an entire paper. Marshall et al. [2015, 2016] train an SVM model with documents annotated for risk of bias, creating RobotReviewer system to predict risk of bias levels. Wang et al. [2022b] utilized a BERT-based pipeline for risk of bias prediction. However, several other works find lower agreement rates between LLMs and human annotators Lai et al. [2024], Hasan et al. [2024], Barsby et al. [2024].
Other LLM Extraction Efforts Recent years have seen LLMs benchmarked in extracting "ground truth" from systematic reviews, particularly in the context of data extraction Sun et al. [2024], Gartlehner et al. [2023], multi-document summaries Wang et al. [2022a], Wallace et al. [2020] and relevance prediction Al-Hussaini et al. [2022]. Another use case has been the extraction of PICOs (Participants, Interventions, Comparisons, Outcomes) from studies. While earlier methods achieved strong results Zafar et al. [2023], Hu et al. [2023], Wallace et al. [2016], Dhrangadhariya and Müller [2023], Jin and Szolovits [2020], Kang et al. [2019], a body of recent work has focused on extracting PICOs with LLMs. Fine-tuning Wadhwa et al. [2023] and prompt tuning Tang et al. [2024] have been used to enhance extraction.
Evaluation of Zero-shot LLM Extraction Kartchner et al. [2023] show that ChatGPT performed inconsistently on various meta-analysis extraction tasks. For instance, while it correctly extracted cancer type roughly 90% of the time, treatment type accuracy fell below 25%. This inconsistency is a considerable motive to more comprehensively understand LLM reliability on various meta-analysis extraction tasks.
Limitation Tasks in Scientific Documents Early efforts in assessing papers’ limitations focused on performing summary extraction and evaluating the result using human evaluation Cohan et al. [2018, 2022], Liu and Shah [2023].
For understanding scientific documents, identifying and retrieving a paper’s limitation is essential to understand the utility and viability of claims and evidences presented in the paper. Faizullah et al. [2024] evaluate fine-tuned embedding models and LLMs’ ability to generate limitation using the extracted limitation from the original paper.
Marshall et al. [2016] use direct quotes in risk of bias judgment to obtain annotations for review authors’ judgments, where exact matching texts are labeled as relevant to a risk of bias. Suster et al. [2023] present a quality assessment task to assess a biomedical paper using the Grading of Recommendation, Assessment, Development, and Evaluation with justifications.
7 Conclusion
We have presented RoBBR, a benchmark for assessing the ability to judge risk of bias in biomedical papers. The benchmark is based on gold-standard risk of bias assessments from Cochrane, and uses a human-validated annotation pipeline for scalably constructing task examples. The four benchmark tasks evaluate different skills required for these assessments, including retrieval and reasoning.
Performance of state-of-the-art language models on the benchmark tasks reveals that current models fall significantly short of expert-level performance. By providing a standardized tool for measuring risk of bias judgments, the RoBBR benchmark can help guide the development of AI systems that aim to automatically assess study quality and synthesize scientific evidence. As systems that review the scientific literature are developed, it is important that they can reliably judge the strength of evidence in order to draw sound conclusions. ML systems which perform well on this benchmark could potentially reduce the time required for risk of bias assessments in systematic reviews, which currently averages nearly six hours per study Crocker et al. [2023]. We hope the RoBBR benchmark will spur further research into developing trustworthy systems for this task.
8 Limitations
The RoBBR benchmark has several limitations. The source of all bias judgments in the benchmark is Cochrane meta-analyses. While these are considered very high quality meta-analyses among biomedical researchers, the bias judgments that they contain are likely to have occasional errors. In addition, there can be disagreements in judgment among different meta-analysis authors.
The Cochrane risk of bias framework does not necessarily identify all types of bias in research studies. The framework uses a predefined set of bias categories, which may not be appropriate for all studies.
The benchmark only includes interventional studies, and therefore excludes observational studies. Assessment of risk of bias in observational studies requires a different procedure than for interventional studies. Our benchmark does not measure performance for observational studies.
Only a limited number of prompting strategies were evaluated for the LLMs. It is likely that improved prompting and task presentation would improve performance on the tasks.
Appendix: Table of Contents
  1. A.
    DatasetA
    1. 1.
      Dataset License and Code LicenseA.1
    2. 2.
      Dataset Hosting, Accessibility and MaintenanceA.2
    3. 3.
      Dataset StatisticsA.3
    4. 4.
      Dataset Collection and Processing A.4
    5. 5.
      RoBBR Structure A.5
  2. B.
    Annotation GuidelineB
  3. C.
    Task 2 Bias Retrieval Prompt and OptimizationC
    • Aspect Decomposition Prompt C.1
    • Aspect Filtering Prompt C.2
    • Aspect-to-Sentence Mapping Prompt and Optimization C.3
    • A Motivating Example C.4
  4. D.
    Task 3 Synthetic Option Generation Procedure and PromptD
  5. E.
    Experiment DetailsE
    • Prompt Optimization E.1
    • Prompts for Evaluation E.2
      • Prompt to Evaluate Study Inclusion/Exclusion (Task 1) E.2.1
      • Prompt to Evaluate Bias Retrieval (Task 2) E.2.2
      • Prompt to Evaluate Support Judgment Selection (Task 3) E.2.3
      • Prompt to Evaluate Risk Level Determination (Task 4) E.2.4
    • Results with Standard Error E.3
  6. F.
    Full Example F
    • Search Protocol F.1
    • Objective of the Meta-Analysis F.2
    • Full Paper F.3
    • Bias Name and Definition F.4
    • PICO F.5
Appendix A Dataset
A.1 Dataset License and Code License
The RoBBR dataset is made available under CC-BY-NC. A copy of the full license can be found at https://github.com/RoBBR-Benchmark/RoBBR/blob/main/LICENSE.md.
The code used in this paper is released under the MIT License. The MIT License is a permissive open-source license that allows for the free use, modification, and distribution of the code, as long as the original license is included with any derivative work. A copy of the full license can be found at https://github.com/RoBBR-Benchmark/RoBBR/blob/main/LICENSE.md.
A.2 Dataset Hosting, Accessibility and Maintenance
The RoBBR dataset with its meta-data is released and can be accessed freely at (https://github.com/RoBBR-Benchmark/RoBBR/blob/main/LICENSE.md). We commit to regularly maintain the dataset and codebase by incorporating user feedback. We will potentially introduce more features as part of future work in the next version of RoBBR. We confirm that the current version of RoBBR will always remain accessible at the same link.
A.3 Dataset Statistics
See table 6 7 8 9 for detailed test set statistics of the four tasks.
Table 6: Test set statistics for Task 1 (Study Inclusion/Exclusion)
# of Tokens in Query Correct Option Count
Dataset Points min avg max Included Excluded
Test Set 309 2685 9380 28464 128 181
Table 7: Test set statistics for Task 2 (Risk of Bias: Sentence Retrieval)
# of Tokens in the Query Sentences Covered Aspects Optimal # of sentences
Dataset Points min avg max min avg max min avg max min avg max
Test Set 513 4104 11220 21494 87 220.1 400 1 1.9 10 1 1.4 9
Table 8: Test set statistics Task 3 (Risk of Bias: Support Judgement Selection)
# of Tokens in the Query
Dataset Points min avg max
Test Set 479 3515 10583 21187
Table 9: Test set statistics for Task 4 (Risk of Bias: Risk Level Determination)
# of Tokens in Query Correct Option Count
Dataset Points min avg max Low Unclear High
Test Set 1052 3256 9763 19834 676 213 163
Table 10: Test set category statistics for Task 2, 3, and 4
Bias Type
Task Points Selection Performance Detection Reporting Attrition Other Analysis Paper specific
Task 2 513 202 81 82 35 95 23 13 7
Task 3 479 130 87 82 80 98 0 8 6
Task 4 1052 332 170 168 119 148 102 18 26
A.4 Dataset Collection and Processing
[…]
A.5 RoBBR Structure
  • task1_SIE_test.json and task1_SIE_dev.json: Test and dev set of task 1 Study Inclusion/Exclusion
    • paper_doi: The DOI of the paper.
    • objective: The meta-analysis objective.
    • search_protocol: search protocol information of the meta-analysis.
    • full_paper: The full paper content.
    • label: One of [included, excluded], showing whether the paper is included or excluded in the meta-analysis.
  • task2_ROBSR_test.json and task2_ROBSR_dev.json: Test and dev set of task 2 Risk of Bias Sentence Retrieval
    • paper_doi: The DOI of the paper.
    • bias: The bias to be considered.
    • PICO: PICO of a study in the paper, including Methods, Participants, Intervention, Outcome, and Notes.
    • objective: The meta-analysis objective.
    • paper_as_candidate_pool: A tuple of text elements from the paper. Each text element is a sentence, a section title, a table, or a figure caption.
    • aspects: A dictionary that maps aspect id to bias aspect.
    • aspect2sentence_indices: a mapping (i.e. dictionary) between aspect id and all sentence indices that independently are source of information for that aspect, as annotated by our pipeline.
    • sentence_index2aspects: a mapping (i.e. dictionary) between sentence index and all aspect ids that this sentence is the source of information of.
    • bias_retrieval_at_optimal_evaluation: This is a dictionary that contains the necessary information for evaluating the model’s performance on the task Bias Retrieval @Optimal.
        optimal: A positive integer, which is the smallest number of sentences needed to cover the largest number of aspects.
        one_selection_of_sentences: a list of sentence indices. The list size is the optimal number. The list of sentences cover the largest number of aspects. Note, there are potentially other lists of sentences that has the same size and also cover the largest number of aspects.
        covered_aspects: the list of aspects that are covered. In this case, the list of aspects covered is list of all aspects.
    • bias_retrieval_at_3_evaluation: This is a dictionary that contains the necessary information for evaluating your model’s performance on the task Bias Retrieval @3.
        one_selection_of_sentences: a list of 3 sentence indices. The list of sentences cover the largest number of aspects that can be covered under the restriction of 3 sentences. Note, there are potentially other lists of sentences that has size 3 and cover the same number of aspects.
        covered_aspects: the list of aspects that are covered. In this case, this list of aspects may not be all the aspects. Since in the paper, we calculate aspect recall by diving the number of aspects covered by model’s retrieved sentences against the total number of aspects, for Bias Retrieval @3, the maximum possible performance is not 100%.
  • task3_SJS_test.json and task3_SJS_dev.json: Test and dev set of task 3 Support Judgment Selection
    • paper_doi: The DOI of the paper.
    • bias: The bias to be considered.
    • PICO: PICO of a study in the paper, including Methods, Participants, Intervention, Outcome, and Notes.
    • objective: The meta-analysis objective.
    • full_paper: The full paper content.
    • options: The seven options for the multiple choice.
    • label: The index of the correct option.
  • task4_RLD_test.json and task4_RLD_dev.json: Test and dev set of task 4 Risk Level Determination
    • paper_doi: The DOI of the paper.
    • bias: The bias to be considered.
    • PICO: PICO of a study in the paper, including Methods, Participants, Intervention, Outcome, and Notes.
    • objective: The meta-analysis objective.
    • full_paper: The full paper content.
    • label: One of [low,high,unclear], representing the risk level of the bias.
Appendix B Annotation Guideline
Below, we show the annotation guideline for aspect mapping of 50 randomly sampled (aspect, full paper) pair. The four annotators form two teams of two persons, all seeing the same annotation guideline.
You have a total of 50 annotation task packets. Each task packet is a docx. file that contains the following information.
  • An Aspect (one piece of important information/detail).
  • The support judgment and the bias.
  • The doi of the paper.
  • An indexed list of text elements of the paper (a text element could be a sentence, tables, a figure caption, etc.)
You have to pledge the following conditions are met during annotation for each task packet.
  • For words you are not familiar with and believe are important for comprehension, conduct the search to understand its meaning.
  • For every text element in the list, you have to look at it and read it at least once.
  • You cannot talk to the other annotator team about anything related to your task, including progress and insights.
  • You should independently do the task first, and then consult with the other person in your team after completing the task.
  • You have to take a mandatory 5 minute break after every 1 hour of performing annotation.
  • You cannot exceed 8 hours of annotation per day.
Below is the recommended procedure for annotating each packet.
  • Read and understand the aspect and the support judgment first. You need to understand the context of the aspect, i.e., the role of the aspect in the support judgment given the bias. You also need to understand why this aspect is important for judging the bias. You can consult with LLM to understand this aspect.
  • Decide what details are important in the aspect. Geo, temporal, and numerical data are all important details.
  • For each text element, you need to decide if a significant amount of important details can be implied from the text element.
    • When deciding the level of implication, consider how the text element can help judge the bias. You should not make complicated implications, i.e., can you see the aspect from the text element within 30 seconds? If not, then it is not a match.
    • Pay attention to acronyms, abbreviations, or different presentation formats of the same information.
    • Even if you find significant information that can be implied from the text element, you have to make sure the text element is in the same context as the aspect.
        Same context typically refers to the same study or experiment, and for numerical results in Table, it means row and column must indicate the same setting.
        Check if the text element refers to the same experiment as the aspect. Since different experiments could be in one paper. Maybe the text element does not refer to any experiment at all.
    • If you suspect that there might be a relationship between the text element and the aspect but do not understand the meaning of the text element, you can use LLM for help. However, you must justify the response from LLM, and you should not rely on the response from LLM.
  • If a text element is a Table, refer to the actual Table in the pdf for better understanding. However, only consider the information in the text element.
  • After you have independently finished all 50 tasks, you should talk to the other person in your team following these procedures:
    • Go through task 0-49.
    • Resolve your difference, check if you made a mistake, or if you missed something. If you made a conceptual error (e.g., fail to understand some terminology), you may have to quickly go through the paper again.
    • For sentences that you cannot resolve your difference after discussion, i.e., one person says yes and the other person says no, or if both people are unsure, you should include them in your final list of decisions.
  • Ultimately, you and your teammate should collaboratively arrive at a consensus for each of the 50 tasks. Write the collective answers in the answer file provided.
Appendix C Task 2 Bias Retrieval Prompt and Optimization
C.1 Aspect Decomposition Prompt
See Figure 6 for the aspect decomposition prompt.
Figure 6: Prompt for decomposing support judgment into aspects.
C.2 Aspect Filtering Prompt
See Figure 7 for prompt to filter negations.
See Figure 8 for prompt to filter non-specific commentaries.
Figure 7: Prompt for extracting negations.
Figure 8: Prompt for filtering non-specific commentaries.
C.3 Aspect-to-Sentence Mapping Prompt and Optimization
To enhance the agreement between human annotators and the GPT4-0125-turbo model, we developed and tested various prompt engineering strategies on a development set consisting of 30 unique (aspect, paper) pairs. It is important to note that no paper or aspect from the development set overlapped with the 50 tasks in the hypothesis testing set.
GPT4-0125-preview was chosen due to its robust instruction-following capabilities and our familiarity with its performance across different settings. Our experimentation revealed that using in-context learning examples did not improve agreement rates and tended to cause overfitting. Instead, we found that embedding the same instruction at both the beginning and the end of the prompt effectively helped the model maintain focus on the task of identifying text elements that significantly cover the details specified in an aspect.
We also implemented chain-of-thought reasoning, prompting GPT4 to articulate its thought process following a specific keyword. This approach not only enhanced the quality of the model’s reasoning but also stabilized its performance and reduced common-sense errors.
To address the issue of entity forgetting in long-context tasks (a typical paper might contain around 5000 tokens), we employed a sliding window technique. Each window, containing 10 text elements with a 5-element overlap, allowed GPT4 to process and evaluate each text element within a manageable context size. The hyperparameters, window length and overlap size, were optimized using the development set. This overlapping approach ensures that each text element is evaluated twice, significantly reducing the likelihood of false negatives.
Figure 9 illustrates the final prompt template used to match text elements (such as sentences, tables, figure captions, etc.) with the aspects. Each template inputs one aspect, the 10 text elements within a sliding window, and the contextual background (specifically, the summary of the IARC paper from which the aspect was derived).
This methodology not only ensures high fidelity in aspect-text alignment but also leverages the model’s capabilities to provide consistent and accurate annotations across extensive text bodies.
Figure 9: Prompt for generating synthetic options.
C.4 A Motivating Example
Here we provide an example of how we build the bias retrieval task.
Bias
Selective reporting (reporting bias)
Support Judgment For the Bias
The trial authors’ original plan was to use Indirect Fluorescent Antibody Test (IFAT) at 6 and 12 months, but due to operational problems, data on IFAT results were not considered valid for the analysis, and serology was not used as a marker of infection in the trial. Problems with serology were poor sensitivity and reproducibility. The authors decided not to use IFAT results in the trial and relied on conversion of the Montenegro skin test (MST) at 18 months of follow-up as the only outcome measure, since no clinical cases of Visceral Leishmaniasis (VL) were detected among the studied population.
Decomposition of Support Judgment into Aspects
  • Aspect 1: The trial authors’ original plan was to use IFAT test at 6 and 12 months
  • Aspect 2: Due to operational problems, data on IFAT results were not considered valid for the analysis
  • Aspect 3: Serology was not used as a marker of infection in the trial
  • Aspect 4: Problems with serology were poor sensitivity and reproducibility.
  • Aspect 5: The authors decided not to use Indirect Fluorescent Antibody Test (IFAT) results in the trial
  • Aspect 6: The authors relied on conversion of the Montenegro skin test (MST) at 18 months of follow-up as the only outcome measure
  • Aspect 7: No clinical cases of VL were detected among the studied population
Aspect Filtering
Using prompt 8, Aspect 4 is filtered since it is a commentary of the reviewer.
Mapping Aspects to Sentences in the paper
Utilizing the procedure described in section C.3, we map the remaining 6 aspects to all sentences in the paper Werneck et al. [2014].
We only show sentences from the paper that are matched with at least one aspect.
  • Sentence 4: The main outcome is the incidence of infection assessed by the conversion of the Montenegro skin test (MST) after 18 months of follow-up in residents aged $\geq$1 year with no previous history of visceral leishmaniasis (VL). (Mapped with aspect 6)
  • Sentence 66: The original plan was to repeat the IFAT test at 6 and 12 months, but due to operational problems, data on IFAT results were not considered valid for the analysis, and serology was not used as a marker of infection in the study. (Mapped with aspect 2, 3, 5, 7)
  • Sentence 73: In any case, we decided not to use IFAT results in this study and relied on conversion of the MST at 18 months of follow-up as the only outcome measure, since no clinical cases of VL were detected among the studied population. (Mapped with aspect 1, 3, 5, 6)
Bias Recall @ Optimal
One of our evaluation metric, Bias Recall @ Optimal, measures the the optimal number of sentences to cover all aspects. In this example, we only need two sentences, sentence 66 and sentence 73, to cover all aspects.
Appendix D Task 3 Synthetic Option Generation Procedure and Prompt
We first generate three detailed synthetic options that are specifically crafted to imitate the support judgments from other papers concerning the same type of bias, ensuring that they are tailored to be relevant to the specific paper in question while maintaining the foundational reasoning. See Figure 10 for the prompt template. Following this, we condense these detailed options into shorter versions that preserve the original meaning, using prompt 11. To prevent heuristic algorithms from solving the task easily, we randomly select either the long or short version of each synthetic option to include in the multiple choice questions.
For selecting the three options derived from other papers’ support judgments within the same bias category, we use prompt 12.
Finally, once the six incorrect options are constructed, we conduct a manual review with the help of prompt 13 on all data points. This review ensures that the options are incorrect and not false negatives.
Figure 10: Prompt for generating synthetic options.
Figure 11: Prompt for shortening the synthetic option.
Figure 12: Prompt for finding negative options from other papers’ support judgments of the same bias.
Figure 13: Prompt for filtering false negatives.
Appendix E Experiment Details
Due to budget limit, we only evaluate the LLMs on a subset of our test set. We provide this subset in our GitHub.
E.1 Prompt Optimization
All evaluation prompts are optimized using disjoint development set for each task. Each prompt includes specialized instructions designed to elicit chain-of-thought reasoning, thereby enhancing the models’ reasoning abilities. These instructions are repeated twice: once at the beginning and once at the end of the prompt, ensuring that the models retain the instructions even after processing the entire paper. Empirical evidence shows that few-shot in-context learning does not improve model performance. Consequently, evaluation prompts do not incorporate in-context learning. It is important to note that prompt optimization is not tailored to any specific large language model (LLM); instead, the aggregated performance across all models is used to guide the optimization of evaluation prompts.
E.2 Prompts for Evaluation
E.2.1 Prompt to Evaluate Study Inclusion/Exclusion (Task 1)
See Figure 14 for prompt to evaluate Study Inclusion/Exclusion (Task 1).
Figure 14: Prompt for Study Inclusion/Exclusion (Task1) evaluation
E.2.2 Prompt to Evaluate Bias Retrieval (Task 2)
See Figure 15 for prompt to evaluate Bias Retrieval (Task 2). We use an additional multi-turn prompt 16 when the model outputs more than required.
Figure 15: Prompt for Bias Retrieval (Task 2) evaluation
Figure 16: Prompt for Bias Retrieval (Task 2) evaluation when the model output more than required
E.2.3 Prompt to Evaluate Support Judgment Selection (Task 3)
See Figure 17 for prompt to evaluate Support Judgment Selection (Task 3).
Figure 17: Prompt for Support Judgment Selection (Task 3) evaluation
E.2.4 Prompt to Evaluate Risk Level Determination (Task 4)
See Figure 18 for prompt to evaluate Risk Level Determination (Task 4).
Figure 18: Prompt for Risk Level Determination (Task 4) evaluation
E.3 Results with Standard Error
We provide bootstrapped standard errors for the experimental results 11,12,13,14.
Table 11: Results for Task 1: Study Inclusion/Exclusion Standard Error Included
F1 Accuracy Weighted Accuracy
Claude3-Opus 75.5 $\pm$ 2.6 77.1 $\pm$ 2.5 74.9 $\pm$ 2.5
Gemini-1.5-Pro 72.2 $\pm$ 2.6 72.7 $\pm$ 2.6 73.0 $\pm$ 2.7
GPT-4o 75.5 $\pm$ 2.6 75.6 $\pm$ 2.5 77.6 $\pm$ 2.3
Table 12: Results for Task 2: Sentence Retrieval Standard Error Included
Bias Type
Model Full Selection Attrition Performance Detection Reporting Analysis Special Other
n = 337 n = 117 n = 57 n= 50 n = 50 n = 35 n = 13 n = 7 n = 23
Bias Recall@Optimal
Claude3-Opus 39.0 $\pm$ 2.4 55.4 $\pm$ 4.0 22.0 $\pm$ 4.7 57.3 $\pm$ 6.39 46.4 $\pm$ 6.5 11.0 $\pm$ 4.9 47.9 $\pm$ 12.3 13.9 $\pm$ 8.3 4.4 $\pm$ 2.9
Gemini-1.5-Pro 41.1 $\pm$ 2.5 57.2 $\pm$ 4.0 28.7 $\pm$ 5.2 49.1 $\pm$ 6.4 48.8 $\pm$ 6.3 23.3 $\pm$ 6.4 54.2 $\pm$ 11.4 13.9 $\pm$ 8.3 3.3 $\pm$ 2.3
GPT-4o 44.8 $\pm$ 2.5 60.3 $\pm$ 4.0 28.6 $\pm$ 5.0 64.6 $\pm$ 5.5 48.7 $\pm$ 6.4 32.9 $\pm$ 7.3 28.7 $\pm$ 9.7 28.0 $\pm$ 14.1 2.3 $\pm$ 2.2
Bias Recall@3
Claude3-Opus 52.3 $\pm$ 2.5 67.9 $\pm$ 3.8 40.2 $\pm$ 5.8 63.1 $\pm$ 5.9 55.9 $\pm$ 6.6 25.5 $\pm$ 7.1 55.9 $\pm$ 12.3 64.6 $\pm$ 16.9 19.7 $\pm$ 7.8
Gemini-1.5-Pro 56.1 $\pm$ 2.4 69.7 $\pm$ 3.8 44.1 $\pm$ 5.7 73.7 $\pm$ 5.3 65.3 $\pm$ 6.1 35.6 $\pm$ 7.6 56.3 $\pm$ 12.8 35.5 $\pm$ 16.1 6.7 $\pm$ 4.8
GPT-4o 61.8 $\pm$ 2.4 72.2 $\pm$ 3.8 47.9 $\pm$ 5.8 82.0 $\pm$ 4.5 78.5 $\pm$ 5.4 43.2 $\pm$ 8.2 67.9 $\pm$ 10.3 50.4 $\pm$ 17.1 10.3 $\pm$ 5.8
Table 13: Results for Task 3: Support Judgment Selection Standard Error Included
Bias Type
Model Full Selection Attrition Performance Detection Reporting Analysis Special
n = 310 n = 86 n = 60 n = 52 n = 56 n = 50 n = 8 n = 6
Accuracy
Claude3-Opus 51.2 $\pm$ 2.8 62.6 $\pm$ 5.2 56.7 $\pm$ 6.4 48.0 $\pm$ 6.9 46.5 $\pm$ 6.9 44.2 $\pm$ 7.0 50.0 $\pm$ 17.9 0.0 $\pm$ 0.0
Gemini-1.5-Pro 49.3 $\pm$ 3.0 57.1 $\pm$ 5.3 59.9 $\pm$ 6.5 46.4 $\pm$ 7.0 42.8 $\pm$ 6.8 40.2 $\pm$ 7.0 49.3 $\pm$ 17.7 16.6 $\pm$ 15.5
GPT-4o 45.8 $\pm$ 2.9 52.3 $\pm$ 5.4 63.1 $\pm$ 6.2 42.5 $\pm$ 6.7 37.7 $\pm$ 6.5 28.0 $\pm$ 6.2 74.5 $\pm$ 15.7 0.0 $\pm$ 0.0
Table 14: Results for Task 4: Risk Level Determination Standard Error Included
Bias Type
Model Full Selection Attrition Performance Detection Reporting Analysis Special Other
n = 337 n = 93 n = 50 n = 50 n = 50 n = 50 n = 18 n = 26 n = 50
F1
Claude3-Opus 47.7 $\pm$ 2.9 54.3 $\pm$ 6.6 34.1 $\pm$ 7.0 46.7 $\pm$ 6.8 37.4 $\pm$ 7.6 23.0 $\pm$ 4.7 69.9 $\pm$ 18.8 27.9 $\pm$ 6.7 27.9 $\pm$ 1.8
Gemini-1.5-Pro 43.1 $\pm$ 2.6 53.9 $\pm$ 5.8 26.0 $\pm$ 3.7 43.7 $\pm$ 7.2 46.8 $\pm$ 7.4 30.4 $\pm$ 5.2 44.6 $\pm$ 7.7 20.1 $\pm$ 3.4 22.8 $\pm$ 6.1
GPT-4o 47.4 $\pm$ 2.9 54.3 $\pm$ 6.6 30.4 $\pm$ 5.8 46.7 $\pm$ 6.8 37.4 $\pm$ 7.6 22.6 $\pm$ 3.9 69.9 $\pm$ 18.8 32.2 $\pm$ 8.5 27.0 $\pm$ 1.6
Accuracy
Claude3-Opus 56.8 $\pm$ 2.6 64.6 $\pm$ 5.0 58.1 $\pm$ 7.1 48.2 $\pm$ 7.1 41.7 $\pm$ 7.1 40.0 $\pm$ 7.3 88.9 $\pm$ 7.3 53.8 $\pm$ 9.7 72.2 $\pm$ 6.4
Gemini-1.5-Pro 49.3 $\pm$ 2.7 62.5 $\pm$ 5.1 42.4 $\pm$ 7.2 46.5 $\pm$ 6.8 48.1 $\pm$ 7.1 56.0 $\pm$ 7.2 77.5 $\pm$ 9.5 42.4 $\pm$ 9.9 23.7 $\pm$ 6.2
GPT-4o 57.0 $\pm$ 2.6 64.6 $\pm$ 5.0 58.1 $\pm$ 7.2 48.2 $\pm$ 7.1 41.7 $\pm$ 7.2 43.6 $\pm$ 7.2 88.9 $\pm$ 7.3 57.7 $\pm$ 9.7 68.3 $\pm$ 6.5
Weighted Accuracy
Claude3-Opus 49.5 $\pm$ 3.0 55.3 $\pm$ 5.1 36.2 $\pm$ 9.0 48.0 $\pm$ 7.0 46.5 $\pm$ 5.9 30.4 $\pm$ 10.1 69.3 $\pm$ 19.6 28.2 $\pm$ 4.9 32.4 $\pm$ 0.9
Gemini-1.5-Pro 47.9 $\pm$ 3.2 56.3 $\pm$ 5.6 34.0 $\pm$ 10.1 44.4 $\pm$ 7.0 54.2 $\pm$ 8.2 36.4 $\pm$ 10.1 43.6 $\pm$ 4.2 22.2 $\pm$ 4.9 32.5 $\pm$ 10.1
GPT-4o 49.3 $\pm$ 3.1 55.3 $\pm$ 5.1 29.3 $\pm$ 3.9 48.0 $\pm$ 7.0 46.5 $\pm$ 5.9 24.5 $\pm$ 8.4 69.3 $\pm$ 19.6 60.4 $\pm$ 24.4 30.6 $\pm$ 1.5
Appendix F Full Example
We provide the full version of the example presented in the main paper. The example is shown in its entirety here, whereas in the main paper it was shortened to save space.
F.1 Search Protocol Gonzalez et al. [2015]
Types of Studies Randomized controlled trials (RCTs)
Types of Participants People living in leishmaniasis endemic regions.
Types of Interventions Any intervention that aims to reduce leishmaniasis incidence through vector or reservoir control.
Types of Outcome Measures People developing CL or VL infections.
Estimates of the vector density measured by an appropriate technique (adult sandfly density estimated by counts of vectors either landing on exposed body parts of humans acting as baits or collected resting inside buildings, for example, on walls).
Number of participants with positive immunological or biochemical tests that detect contact with the parasite (for example, leishmanin skin test conversion rates or lymphocyte proliferation rates, or both).
Adverse effects on people.
Adherence to control measures; for example, the extent to which specified intervention components were delivered as prescribed.
Measures of environmental impact (assessment of the possible impact - positive or negative - that the interventions may have on the natural environment) or sustainability (assessment of the ability to change biological and human processes, functions, biodiversity and productivity), or both.
F.2 Objective of the Meta-Analysis Gonzalez et al. [2015]
To assess the effects of vector and reservoir control interventions for cutaneous and for visceral leishmaniasis.
F.3 Full Paper
See Werneck et al. for full paper content.
F.4 Bias Name and Definition Higgins and Green [2011]
Bias Name
selective reporting (reporting bias)
Bias Definition
Reporting bias due to selective outcome reporting.
F.5 PICO
Methods Trial design: Cluster-RCT. Unit of randomization: Geographic area. Number of clusters: 40 geographic areas. Entomological data collection: Not done. Clinical data collection: Conversion of the Montenegro skin test (MST) at 18 months of follow-up. Length of follow-up: 18 months. Analysis: Analysed at cluster level.
Participants Ten localities in 7 neighbourhoods of the city of Teresina (Brazil) were divided into blocks, each containing an average of 60 residences. For each locality, 4 blocks were selected to minimize the risk of cross-contamination of interventions. Eligible participants were residents of selected blocks aged 1 year or above with no history of VL. The 40 geographic areas (blocks) randomly allocated to the 4 types of interventions (697 subjects MST-). Endemic disease: VL caused by L. chagasi (L. infantum).
Interventions 1. Spraying households and residential annexes with insecticide. 2. Elimination of infected dogs. 3. Combination of spraying and eliminating infected dogs. 4. No intervention. Description of spraying: performed according to the routine of the VL Control Program of the Zoonosis Control Center of the Teresina City Health Department. Interventions were delivered in the selected blocks every 6 months, for three times, beginning just after each household visit. The elimination of infected dogs was decided if indirect immunofluorescence test was more or equalled 1:40.
Outcomes Cases of infection by L. infantum at 18 months determined by conversion of the MST (MST- at the beginning) or diagnosis of active VL.
Notes Country: Brazil (Teresina, Itararé quarter). Trial dates: January 2004 to December 2006. Trial sponsor: Funded by Health Surveillance Unit from the Brazilian Ministry of Health. One author was partially funded by the Brazilian Research Council (CNPq 306267/2010-1 and 202088/2012-0). The founders had no role in trial design, data collection and analysis, decision to publish, or preparation of the manuscript. The authors have declared that no competing interests exist. Sample size: Calculated. Compliance assessment: Not reported.
[…]

### v2

@@
-supSupplementary References
+Measuring Risk of Bias in Biomedical Reports: The RoBBR Benchmark
 Abstract
 Systems that answer questions by reviewing the scientific literature are becoming increasingly feasible.
-To draw reliable conclusions, these systems should take into account the quality of available evidence, placing more weight on studies that use a valid methodology.
+To draw reliable conclusions, these systems should take into account the quality of available evidence from different studies, placing more weight on studies that use a valid methodology.
 We present a benchmark for measuring the methodological strength of biomedical papers, drawing on the risk-of-bias framework used for systematic reviews.
-The four benchmark tasks, drawn from more than 500 papers, cover the analysis of research study methodology, followed by evaluation of risk of bias in these studies.
-The benchmark contains 2000 expert-generated bias annotations, and a human-validated pipeline for fine-grained alignment with research paper content.
-We evaluate a range of large language models on the benchmark, and find that these models fall significantly short of expert-level performance.
-By providing a standardized tool for measuring judgments of study quality, the benchmark can help to guide systems that perform large-scale aggregation of scientific data.
+Derived from over 500 biomedical studies, the three benchmark tasks encompass expert reviewers’ judgments of studies’ research methodologies, including the assessments of risk of bias within these studies.
+The benchmark contains a human-validated annotation pipeline for fine-grained alignment of reviewers’ judgments with research paper sentences.
+Our analyses show that large language models’ reasoning and retrieval capabilities impact their effectiveness with risk-of-bias assessment.
 The dataset is available at https://github.com/RoBBR-Benchmark/RoBBR.
+Measuring Risk of Bias in Biomedical Reports: The RoBBR Benchmark
+Jianyou Wangthanks: Equal contribution   Weili Cao11footnotemark: 1   Longtian Bao   Youze Zheng   Gil Pasternak Kaicheng WangXiaoyue WangRamamohan PaturiLeon Bergen Laboratory for Emerging Intelligence University of California, San Diego {jiw101, w2cao, rpaturi, lbergen}@ucsd.edu
+Figure 1: Reviewer finds Sentence 4, 66, 73, 183.
+The reviewer writes a support judgment based on these sentences.
+The support judgment indicates this biomedical study has a high risk for selective reporting bias.
+Both support judgment and the high risk rating are included in the systematic review.
 1 Introduction
 Systems that automatically answer questions by reviewing the scientific literature are becoming increasingly feasible.
-These systems offer the potential to provide scientists with on-demand access to knowledge that is synthesized from across the literature.
+These systems, such as "Deep Research" offerings from OpenAI, Google, and Anthropic, have the potential to provide scientists with on-demand access to knowledge that is synthesized from across the literature.
 This can help scientists understand what is already known about topics outside of their research focus, and help clinicians and practitioners keep up to date with best practices.
-When assessing what is known about a field, not all studies should be weighed equally I et al.
-[2023].
+When assessing what is known about a field, not all studies should be weighed equally Boutron et al.
+(2023).
 Studies with stronger methodologies contribute more to a body of evidence than those with weaker methodologies.
 By weighing studies appropriately, systems can increase the reliability of their summaries and recommendations Turner et al.
-[2009].
-In biomedical research, there is a large body of work which investigates the factors that decrease the validity of a study’s methods, and best practices for addressing these issues I et al.
-[2023], Welton et al.
-[2009].
-Reporting bias occurs when there are systematic differences in how outcomes are reported or disclosed between the groups that are compared (e.g., the results of the treatment group are more frequently or favorably published than those in the placebo group).
-This bias is often best addressed by adhering to strict reporting standards and protocols, such as registering the study in advance and committing to publish all results Kotz and West [2022].
-Other biases, such as detection bias, attrition bias, and selection biases, can frequently be addressed using careful study design and implementation Berger [2005], JPT et al.
-[2023], I et al.
-[2023].
-This paper presents the RoBBR benchmark for assessing the methodological strength of biomedical studies.
-The benchmark follows the risk of bias framework developed by Cochrane Jpt et al.
-[2023], JPT et al.
-[2023], Higgins et al.
-[2011], Practice and of Care [EPOC], Sterne et al.
-[2014], a global independent network that conducts systematic reviews of healthcare interventions.
-Cochrane’s risk of bias framework is widely recognized as a standard for evaluating the quality and reliability of research studies.
-It serves as an important resource for national health agencies in various countries, including the United States Viswanathan et al.
-[2012], the United Kingdom Alderson and Tan [2011], and Australia NHMRC [2019].
-Notably, Cochrane’s systematic reviews are used to formulate national and international guidelines on healthcare practices Bunn et al.
-[2015].
-Furthermore, empirical comparisons between Cochrane and non-Cochrane systematic reviews consistently demonstrate higher methodological quality in those conducted by Cochrane, highlighting its critical role in advancing evidence-based healthcare Goldkuhle et al.
-[2018], Page et al.
-[2018], Matthias et al.
-[2020], Tsoi et al.
-[2020].
-Figure 1: The figure illustrates the components of a risk of bias assessment.
-Given a bias category and a research study, a meta-reviewer identifies factors that make the study high or low risk for that bias.
-The meta-reviewer summarizes their findings in a support judgment, and then make a final decision about the risk level.
+(2009).
+In biomedical research, there is a large body of work which investigates the factors that decrease the validity of a study’s methods, and best practices for addressing these issues Boutron et al.
+(2023); Welton et al.
+(2009).
+For instance, reporting bias occurs when there are systematic differences in how outcomes are reported or disclosed between the groups that are compared (e.g., the results of the treatment group are more frequently or favorably published than those in the placebo group).
+This bias can be mitigated by registering the study in advance and committing to publish all results Kotz and West (2022).
+We introduce the RoBBR benchmark for evaluating the methodological strength of biomedical studies, which is referred to as their risk-of-bias levels in the context of this paper.
+RoBBR’s main task involves labeling these risk-of-bias levels.
+Importantly, we also introduce two novel subtasks: support sentence retrieval (SSR) and support judgment selection (SJS).
+The SSR subtask is designed to test a model’s ability to identify support sentences — specific sentences within an entire paper that signal potential biases.
+We created the SSR dataset using our annotation pipeline.
+The SJS subtask evaluates a model’s capacity to synthesize this retrieved information through reasoning to form a well-supported judgment.
+Through our in-depth analysis, we find that the abilities to reason and, especially, to retrieve information are important milestones in improving models’ capacity to assess risk-of-bias levels.
 2 Background
-The RoBBR benchmark aims to provide an expert-validated tool for assessing whether models can make risk of bias judgments for biomedical papers.
-The benchmark uses a risk-of-bias framework developed in the context of systematic reviews, which is the gold standard for assessing biomedical studies.
-We provide more background on systematic reviews and the risk of bias framework below.
-Systematic reviews and meta-analyses are tools used in evidence-based medicine, which synthesize evidence from multiple studies to answer important clinical questions.
-They help researchers and healthcare professionals make informed decisions based on the best available evidence Ahn and Kang [2018], TJ et al.
-[2023], JJ et al.
-[2023].
-A systematic review is a comprehensive search of the literature to identify all relevant studies on a particular topic TJ et al.
-[2023], C et al.
-[2023].
-The review process follows a pre-specified protocol to minimize bias and ensure transparency.
-The protocol outlines the research question, search strategy, study selection criteria, data extraction methods, and plans for quality assessment and synthesis of result Jpt et al.
-[2023], Page et al.
-[2021].
-A meta-analysis is a type of systematic review that uses statistical methods to combine the results from multiple studies Ahn and Kang [2018].
-By pooling the data across studies, a meta-analysis can provide a more precise estimate of the treatment effect JJ et al.
-[2023], Guevara et al.
-[2004] and explore potential sources of heterogeneity Higgins and Thompson [2002].
-The research question for a systematic review is often structured using the PICO framework, which stands for Participant, Intervention, Comparison, and Outcome Aslam and Emmanuel [2010].
-The PICO elements help define the key components of the research question and guide the study selection process J et al.
-[2023], JE et al.
-[2023a].
-For example, a systematic review on the effectiveness of a new hypertension drug might have the following PICO: Participant: Adults with primary hypertension; Intervention: New antihypertensive drug; Comparison: Placebo or standard care; Outcome: Change in systolic and diastolic blood pressure.
-By clearly defining the PICO elements, the review authors can develop a focused search strategy and establish explicit criteria for selecting studies to include in the review J et al.
-[2023], JE et al.
-[2023b].
-Inclusion criteria typically specify the characteristics of the study population, the interventions and comparisons of interest, the outcomes to be measured, and the study designs to be included JE et al.
-[2023b], Page et al.
-[2021], Tawfik et al.
-[2019] (e.g., randomized controlled trials).
-Exclusion criteria describe reasons why a study must be excluded from an analysis, such as using the wrong study design C et al.
-[2023], Page et al.
-[2021], Tawfik et al.
-[2019].
-Risk of bias assessment involves evaluating each included study to determine the likelihood that its results may be biased JPT et al.
-[2023], Practice and of Care [EPOC], Sterne et al.
-[2014].
-The Cochrane tool for assessing risk of bias Higgins et al.
-[2011] is widely used and assesses several types of bias, including random sequence generation, allocation concealment, blinding of participants and outcome assessors, incomplete outcome data, selective reporting, and other sources of bias.
-During the assessment, two review authors independently evaluate the risk of bias using the Cochrane tool and resolve disagreements through discussion.
+We provide more background on systematic reviews and the risk-of-bias guidelines.
+A systematic review is a method in evidence-based medicine that synthesizes evidence from multiple studies to answer important clinical questions.
+They help researchers and healthcare professionals make informed decisions based on the best available evidence Ahn and Kang (2018); Lasserson et al.
+(2023); Deeks et al.
+(2023).
+Risk-of-bias assessment involves evaluating each study to determine the likelihood that its results may be biased Higgins et al.
+(2023); Practice and of Care (EPOC); Sterne et al.
+(2014).
+Studies with a high risk of bias may overestimate or underestimate the true treatment effect, leading to inaccurate conclusions in the systematic review.
+RoBBR follows the risk-of-bias guideline developed by Cochrane Sterne et al.
+(2019); Higgins et al.
+(2023, 2011); Practice and of Care (EPOC); Sterne et al.
+(2014), a global independent network that conducts systematic reviews of healthcare interventions.
+Cochrane’s risk-of-bias guideline is widely recognized as a standard for evaluating the quality and reliability of research studies.
+It serves as an important resource for national health agencies in various countries, and are used to formulate national and international guidelines on healthcare practices Viswanathan et al.
+(2012); Alderson and Tan (2011); NHMRC (2019); Bunn et al.
+(2015).
+During reviewer’s assessment of a study’s risk of bias, two review authors independently evaluate the risk of bias using the Cochrane tool and resolve disagreements through discussion.
 They record their rationale for each bias assessment.
-These rationales are referred to as "support judgments."
-For example, a support judgment of Reporting Bias could be "The study protocol was registered (ClinicalTrials.gov NCT01863706) but not all of the outcomes projected by methodological descriptions were reported as results in the study report (cases of diarrhoea, nausea and vomiting were not completely reported).
-Moreover, the study publication reports outcomes (hypotension, nausea, transfusion) not listed in the registered protocol." A decision regarding the risk of reporting bias is then made on the basis of the support judgment.
-See Figure 1 for an illustration.
-Studies with a high risk of bias may overestimate or underestimate the true treatment effect, leading to inaccurate conclusions in the systematic review.
-In this paper, we consider eight categories of biases: selection bias, attrition bias, performance bias, detection bias, reporting bias, analysis bias, special bias (specific to individual meta-analyses), and other bias (i.e.
-biases not covered by the other categories).
-We benchmark model performance on assessing papers for each of these bias categories.
+These rationales are referred to as "support judgments".
+Figure 2: Given a support judgment, our annotation pipeline produces the Aspect-to-Sentence Mapping, a many-to-many relationship.
+The mapping on the right-side shows Sentence 66 covers both Aspect 1 and 2, so if a model retrieves Sentence 66, it should not retrieve Sentence 4 to avoid redundancy.
+3 Related Work
+Risk-of-Bias Labeling
+Previous research focuses on creating datasets and training models to label risk-of-bias levels Lai et al.
+(2024a); Hasan et al.
+(2024); Wang et al.
+(2022b); Marshall et al.
+(2015).
+Traditionally, machine learning models including support vector machines Marshall et al.
+(2014, 2017); Pereira et al.
+(2020), convolutional neural networks Marshall et al.
+(2017); Zhang et al.
+(2016) and logistic regression Millard et al.
+(2016); Marshall et al.
+(2020) are used for label prediction.
+More recently, some work employs transformers Dias et al.
+(2025) and LLMs to determine risk-of-bias levels Lai et al.
+(2024b); Pitre et al.
+(2023); Šuster et al.
+(2024).
+While RoBBR’s main task is similar to existing work, its novel contribution is the introduction of two new subtasks: support sentence retrieval (SSR) and support judgment selection (SJS).
+These subtasks identify crucial intermediate steps in risk-of-bias labeling and respectively measure the information retrieval and reasoning capabilities of LLMs when they assess risk-of-bias.
+Support Sentences in Risk-of-Bias Labeling
+Previous studies Marshall et al.
+(2016); Dias et al.
+(2025) attempt to extract supporting sentences from reviewers’ support judgments using direct quotes Marshall et al.
+(2016) or semantic embeddings Dias et al.
+(2025).
+However, these techniques falter because support judgments are free-form and frequently omit direct quotes, while semantic embeddings are imprecise and incomprehensive (See Table 2).
+This makes sentences extracted by such methods unsuitable for a rigorous benchmark like our Support Sentence Retrieval (SSR) subtask, which tests an LLM’s capability in retrieving support sentences.
+To address this, in Section 4.3, our annotation pipeline aligns all aspects (such as paraphrases, synthesized and summarized information) from support judgment to sentences in biomedical studies, which would more reliably and comprehensively extract support sentences that are used to form the novel SSR subtask.
+LLM Extraction from Systematic Reviews
+In recent years, LLMs have been used to extract data from systematic reviews, such as experiment evidence Sun et al.
+(2024); Gartlehner et al.
+(2023), summaries Wang et al.
+(2022a); Wallace et al.
+(2020), quality of evidence Suster et al.
+(2023) and PICO Tang et al.
+(2024).
+In our work, aligning the human support judgment to various sentences in biomedical studies is a more complex and difficult "data extraction" task, as noted by Dias et al., which necessitates the aspect decomposition and aspect & sentence annotation components in our annotation pipeline.
 4 Benchmark Development
-The RoBBR benchmark is derived from 63 Cochrane meta-analyses and 532 papers referenced in those meta-analyses.
-All of the meta-analyses and papers have a Creative Commons license or are public domain.
-The benchmark is designed to evaluate a model’s ability to judge the risk of bias in biomedical reports.
-The data were collected in April 2024.
-We included all meta-analyses from Cochrane that are in the BioC database, are classified as intervention reviews, have a CC-BY-NC license, and contribute at least one data point for each task.
-We obtained the XML format of the papers included in these meta-analyses either through the BioC API or by manually downloading and converting PDFs to XML using GROBID.
-We then extract task data from the XML of these papers.
-The benchmark is divided into four main tasks (See in Appendix A.3 for detailed statistics):
-  • Study Inclusion/Exclusion (Task 1): Determine whether a study’s objectives and methodology fit a set of requirements.
-  • Bias Retrieval (Task 2): Retrieve sentences from the paper that support a judgment about the study’s risk of bias.
-  • Support Judgment Selection (Task 3): From a list of options, select the judgment that best supports a determination of the study’s risk of bias.
-  • Risk Level Determination (Task 4): Assess the study’s risk of specific biases.
-These four tasks mirror the procedure followed by meta-reviewers.
-Initially, reviewers decide if a study should be included or excluded based on predefined criteria, which reflect the goals of the review (Task 1).
-Included studies will be analyzed for indications of the risk of bias (Task 2).
-Reviewers will then summarize the evidence collected from the study, and analyze its implications (Task 3).
-Finally, they assess the risk level of specific biases for each identified bias (Task 4).
-By structuring the dataset around these tasks, we provide a comprehensive framework for simulating the nuanced process of risk of bias assessment in meta-analysis.
-3.1 Train/Test Split
-We randomly select 23 meta-analyses and their 228 corresponding papers in the dev set. We put the other 40 meta-analyses and their corresponding 304 papers in the test set.
-Figure 2: Study Inclusion/Exclusion. In Task 1, the model sees a search protocol and objective, and is asked to determine whether a study’s methodology fits those criteria.
-3.2 Task 1: Study Inclusion/Exclusion
-For each potentially relevant study to be included in a meta-analysis, the search protocol information and meta-analysis objective are provided as input, and the task is to decide if the study should be included in the meta-analysis.
-The objective is included in order to provide similar context to what is normally available for a meta-reviewer and make it easier to interpret the criteria in the search protocol.
-An example is provided in Figure 2.
-3.3 Task 2: Risk of Bias Sentence Retrieval
-When judging the risk of bias for a specific study, the review authors provide support judgments which justify their risk of bias rating.
-These support judgments are grounded in specific aspects of the study, which the support judgment describes.
-Task 2 aims to test a model’s ability to retrieve the correct source for the support judgment from a paper’s text.
-The support judgments are intentionally omitted in the Bias Retrieval task.
-This exclusion simulates the systematic review process by challenging the model to independently generate these justifications.
-Including them would come close to providing the answer, undermining the purpose of the task.
-See Figure 3 for an illustration.
-Some review authors directly quote sentences from the paper in their support judgment, allowing for straightforward extraction and matching of these sentences with the paper’s text.
-However, most reviewers prefer to paraphrase the sentences or engage in deeper analysis, synthesizing information from multiple parts of the paper.
-To match these judgments with sentences from the paper, we use a three part annotation pipeline.
-First, a support judgment often contains multiple pieces of information.
-We decompose the support judgment into distinct, non-overlapping pieces, each focusing on a specific aspect of the original support judgment.
-Second, we exclude any part of the support judgment containing a negation or non-specific commentary, because these parts of the judgment cannot be matched to any sentence.
-The remaining parts of the support judgments generally contain specific claims about the study.
-Third, these aspects are mapped to specific sentences within the paper, allowing us to pinpoint the exact sources of information that underlie each aspect of the judgment.
-A single aspect could potentially be supported by multiple sentences from the paper, and similarly a single sentence can support multiple aspects.
-Each of these steps, including the aspect decomposition, uses GPT-4 (gpt-4-0125) OpenAI [2023]. The prompts can be found in the Appendix C.
-We manually inspect the 2221 support judgment aspects generated in the second stage, to ensure that all aspects are related to specific information from the paper.
-In the next section, we evaluate annotation quality by comparing to human judgments.
-Figure 3: Risk of Bias: Sentence Retrieval.
-In Task 2, the goal is to retrieve sentences from the paper (the Candidate Pool) which support a risk of bias judgment.
-Performance is evaluated against a Cochrane review support judgment, which has been split into aspects.
-3.3.1 Evaluating LLM Annotations
-To match each aspect from a support judgment with specific sentences in the paper, we need to evaluate each sentence (on average, 200 sentences) against each aspect.
-For a support judgment with 5 aspects, this results in approximately 1,000 decisions.
-Given the scale of our entire dataset, which includes over 500 (bias, study) pairs, this translates to more than 1 million decisions – an impractical task for human annotators to perform manually.
-We hypothesize that determining whether a sentence matches an aspect is straightforward enough for GPT-4 to achieve human-level performance with a specialized prompt.
-Following the protocol introduced by Wang et al.
-[2023], we created a development set of 30 papers.
-A single support judgment aspect was chosen for each paper, and every sentence in the paper was annotated for that aspect, totaling 8789 decisions.
-We optimized the instruction prompt for GPT-4 on this development set.
-Details about the prompt and experimental procedure are provided in the Appendix C.
-Once the prompt was finalized, four annotators were tasked with annotating 50 randomly sampled (aspect, full paper) pairs (all annotators saw the same pairs).
-Every sentence in each paper was annotated, requiring 13575 annotations in total.
-The annotators were divided into two teams, with each team consisting of two annotators.
-Each annotator performed the annotation tasks individually given a pre-defined annotation guideline.
-Annotators from the same team then collaborated to resolve differences and eliminate mistakes.
-Refer to the Appendix B for the full annotation guidelines.
-Each team produced a set of annotation results.
-GPT-4 annotated the same (aspect, paper) pairs.
-We calculated Binary Accuracy, F1-binary, Spearman Correlation, and Kappa Coefficient between the two human teams and between each human team and GPT-4.
-Table 1 shows that the human teams matched each other and GPT-4 more than 99% of the time.
-The other agreement metrics have values in the low-to-mid 70’s, indicating substantial agreement among annotators.
-[^1] These metrics are lower than exact agreement because of class imbalance.
-Positive matches between sentences and aspects are rare.
-Bootstrapped hypothesis tests do not demonstrate any statistical difference between GPT-4 and the human annotators.
-Table 1: Inter-annotator agreement among the two teams of humans and the human teams and GPT-4.
-Metrics Human & Human Average Human & GPT Average p-value
+4.1 Benchmark Statistics and Data Sources
+Cochrane Cochrane Non-Cochrane
+Task Train Test Test
+Retrieval (SSR) n=235 n=313 N/A
+Selection (SJS) n=346 n=465 N/A
+Risk-of-Bias (Main) n=774 n=906 n=2,489
+Table 1: RoBBR Benchmark Statistics.
+There are two prevailing risk-of-bias assessment guidelines, RoB1 Higgins and Green (2011) and the updated RoB2 Sterne et al.
+(2019).
+As of 2025, many recently published and updated systematic reviews still use RoB1, for example, Miao et al.
+(2025); Mehrholz et al.
+(2025).
+Since the RoB2 guidelines do not require support judgments that are necessary for our subtasks, our subtasks only include systematic reviews that follow RoB1.
+For the main task, we include systematic reviews that follow either set of guidelines.
+Since there are hundreds of bias names and definitions, such as "incomplete adverse event reporting (reporting bias)" and "selective outcome reporting (reporting bias)", for clarity of presentation, we manually group them into 6 primary categories: "selection", "attrition", "performance", "detection", "reporting" and "deviation" (unique to RoB2) using two approaches.
+In some cases, bias names already include broader category labels (e.g., "random sequence generation (selection bias)").
+When this happens, we follow these broader category labels.
+When bias names lack such broader category labels, we manually identify an equivalent bias name that has a category label and use this label.
+For example, the bias name "bias arising from the randomization process" is equivalent to "random sequence generation (selection bias)," so we use "selection bias" as its category label.
+A bias name belongs to multiple categories when these categories are explicitly included in the bias name.
+For example, a bias called "blinding for adverse event (performance and detection bias)" is categorized into both the "performance bias" category and the "detection bias" category.
+See Table 12 for examples of how bias names are categorized.
+During evaluation, models would see the actual bias name and definition used in systematic reviews.
+Bias definitions are sourced from systematic reviews themselves and official guidelines like Higgins and Green (2011); Practice and of Care (EPOC); Sterne et al.
+(2014).
+For data diversity, we include both systematic reviews that are published in Cochrane (denoted as Cochrane Train/Test) and that are published elsewhere (denoted as Non-Cochrane Test [^1] Empirically, Llama3 fine-tuned on Cochrane train set also perform well on Non-Cochrane test set, so we did not create a separate Non-Cochrane train set.).
+Non-Cochrane reviews do not have support judgments and cannot form our subtasks.
+RoBBR includes 58 Cochrane reviews that assess 204 papers, and 496 Non-Cochrane reviews that assess 496 papers.
+Appendix A.3 show details on token length, bias and label distributions.
+See Table 1 for statistics of the three tasks in our benchmark.
+4.2 Main Task: Risk-of-Bias Determination
+The main task evaluates a model’s ability to label the risk-of-bias of a biomedical study as high, low, or unclear/some concern.
+The input to the model is exactly everything that a human expert would see, including the biomedical paper, study characteristics (i.e.
+PICO), the objective/topic of the systematic review, one specific bias name and definition, and risk level definition.
+See the top-left panel of Figure 1 for a visualization of input.
+The same input would be used for the two subtasks as described in section 4.4, 4.5.
+To visualize the task, see Figure 3.
+Figure 3: Main Task: Risk-of-Bias Determination. The goal of the main task is to provide an assessment of the paper’s risk of specific biases.
+The decision-making process of risk-of-bias determination is highly complex and involves multiple intermediate steps: retrieve support sentences from the biomedical paper, synthesize evidence from support sentences into a support judgment, and follow the guidelines and definitions of bias and risk level to give a final determination.
+4.3 Automatic Annotation Pipeline
+Support judgments contain a wealth of information, including evidence, data, reasoning, and commentaries written by expert reviewers.
+The goal of support judgment is to explain why reviewers would give this risk-of-bias rating.
+Therefore, support judgments mirror and shed light on the decision-making process of the reviewer.
+It is difficult to find all support sentences in the study that form the basis of a support judgment.
+As shown in Figure 2, support judgments can have many aspects.
+While some aspects are direct quotes from the study, many other aspects are paraphrases from the study, or deeper analyses that synthesize information from multiple parts of the study.
+To reverse engineer a support judgment, we find that no lexicon-based or embedding models can find all support sentences accurately (see Section 4.3.1 and Table 2).
+To address this, our proposed annotation pipeline has two main components: Aspect Decomposition and Aspect & Sentence Annotation.
+See Figure 2 for a visualization.
+GPT-4 (gpt-4-0125) is used by our annotation pipeline.
+All prompts used in our annotation pipeline are in Appendix D.
+Aspect Decomposition:
+To reduce the complexity of the reverse engineering process, we decompose the support judgment into distinct, non-overlapping pieces of information, each focusing on a specific aspect of the original support judgment.
+This is accomplished by careful prompt engineering with GPT-4.
+We manually validated that all decomposed aspects from all support judgments are of high quality.
+Aspect & Sentence Annotation:
+To further reduce complexity and instability introduced by long context input to LLMs, we only ask GPT-4 to determine if one sentence from the study covers one aspect.
+Definition 1.
+A sentence $s_{j}$ in a paper covers an aspect $f_{i}$ if it satisfies these two criteria:
+1: The content of the sentence conveys the majority of the aspect’s information.
+2: Any information about the aspect not directly stated in the sentence can be reasonably inferred from the surrounding context.
+We aggregate these annotations into the Aspect-to-Sentence Mapping: For each (study, support judgment) triplet, a sentence is mapped to several aspects (e.g., 0, 1, 2, or more).
+Most sentences are not mapped to any aspect because they do not cover them.
+A few sentences can even be mapped to more than one aspect because aspect-to-sentence mapping is a many-to-many relationship.
+The total cost of our annotation pipeline via OpenAI API is around \$1,000 with aspect decomposition and initial filtering costing less than \$100. Next, we evaluate the quality of GPT-4 annotation.
+Metrics Human & Human Human & GPT p-value
 Exact Accuracy 99.4 $\pm$ 0.2 99.3 $\pm$ 0.2 0.31
@@
 Spearman’s $\rho$ 73.8 $\pm$ 6.0 71.7 $\pm$ 5.6 0.43
-3.4 Task 3: Risk of Bias Support Judgment Selection
-The previous task evaluates a model’s ability to extract information from biomedical papers to in order to justify assessments of bias.
-However, this approach alone is insufficient for a risk-of-bias assessment.
-The evaluation must also include the model’s ability to reason about the retrieved information.
-For example, a paper might report the loss to follow-up yet fail to provide reasons for missing data, which raises the possibility of attrition bias.
-Similarly, while a paper may state that it implemented double-blinding, expert reviewers could consider such blinding infeasible due to the nature of the trial, potentially introducing performance bias.
-To address this, we propose a multiple-choice task where the model selects the correct support judgment.
-This task includes three synthetically generated hard options, three options derived from other papers’ support judgments for the same bias category, and the Cochrane reviewer’s actual judgment.
-We prompt GPT-4 to generate the three synthetic options by imitating support judgments from other papers concerning the same type of bias.
+Metrics Human & Human Human & Embedding p-value
+Exact Accuracy 99.4 $\pm$ 0.2 98.5 $\pm$ 0.2 <10e3<\!10e{-3}
+F1 Binary 73.4 $\pm$ 6.3 30.5 $\pm$ 8.1 <10e3<\!10e{-3}
+Cohen’s $\kappa$ 73.1 $\pm$ 6.3 29.7 $\pm$ 8.2 <10e3<\!10e{-3}
+Spearman’s $\rho$ 73.8 $\pm$ 6.0 29.8 $\pm$ 8.3 <10e3<\!10e{-3}
+Table 2: Inter-annotator agreement. Top: comparison between two human teams and GPT-4. Bottom: comparison between two human teams and OpenAI-v3-large as a post-hoc analysis.
+4.3.1 Evaluating Annotation Quality
+We hypothesize a sentence matches an aspect is straightforward enough for GPT-4 to achieve human-level performance with a specialized prompt.
+Following the protocol introduced by Wang et al.
+(2023), we optimized the instruction prompt for GPT-4 on a development set (see Appendix D).
+We randomly sampled 50 papers, with each assigned a single aspect.
+Across all 50 papers, this resulted in a total of 13,575 (aspect, sentence) pairs.
+Four graduate student were tasked with annotating these pairs.
+The annotators were divided into two teams, with each team consisting of two annotators.
+The annotators’ training for annotation is described in Appendix B.1.
+Each annotator performed the annotation tasks individually given a pre-defined annotation guideline in Appendix B.
+Annotators from the same team then collaborated to resolve differences and eliminate mistakes.
+Each team produced a set of annotation results.
+GPT-4-0125 annotated the same 13k (aspect, sentence) triplets.
+We calculated Exact Accuracy, F1-binary, Spearman Correlation, and Kappa Coefficient between the two human teams and between each human team and GPT-4.
+Table 2 shows human-human correlations in the low-to-mid 70s, indicating reasonable agreement, despite the fact that finding sentences that cover aspects is non-trivial due to different annotators’ potentially varied interpretations of key terms in Definition 1 such as "majority" and "reasonably inferred".
+For context on inter-annotator agreement in similar aspect-based tasks, Wang et al.
+reported 54% $\rho$ (three-way annotation) and Mysore et al.
+found near 40% pre-adjudication $\rho$ (four-way annotation).
+Table 2 shows human-GPT4 and human-human correlations are comparable.
+A post-hoc analysis reveals that powerful embedding models like OpenAI-v3-large OpenAI (2024) have significantly lower correlation with humans.
+For the primary human-GPT4 comparison, the p-value (p = 0.3) from our bootstrapped test fails to reject the null hypothesis (no difference in correlation).
+In contrast, we can reject the null hypothesis for the embedding model which has clear and detectable differences from human annotators.
+4.4 Support Sentence Retrieval (SSR)
+With our annotation pipeline, we transform a support judgment into a special information retrieval task that evaluates if a model can identify and retrieve these support sentences from the full text of the paper.
+Since these support sentences form the basis of the support judgment, if a model can locate these support sentences, it is one step closer to reaching a good support judgment and classifying the correct risk-of-bias level.
+Figure/table captions are split into sentences as text.
+Tables are included and turned into markdown format and treated as individual sentences.
+Figures are excluded since SSR is designed to evaluate textual models.
+We have obtained the Aspect-to-Sentence Mapping.
+From this mapping, there exists the smallest optimal integer $K$ such that there are $K$ sentences that can cover all aspects in a support judgment.
+Therefore, when a model is only allowed to retrieve $K$ sentences from a study, we measure the percentage of aspects that the model’s retrieved set of sentences can cover.
+This metric is denoted as Aspect Recall Ratio @ Optimal.
+Formally, let $\{s_{1},\dots,s_{k}\}$ be the set of retrieved sentences, and $\{f_{1},\dots,f_{m}\}$ be the set of aspects. We define the indicator function $\mathcal{S}(f_{i},s_{j})$ as:
+$\mathcal{S}(f_{i},s_{j})=\begin{cases}1&\text{if }s_{j}\text{ covers }f_{i}\\ 0&\text{otherwise}\end{cases}$
+The Aspect Recall Ratio @ Optimal is defined as
+$\frac{\sum_{i=1}^{m}\displaystyle\mathbf{1}\Bigg\{{\left(\sum_{j=1}^{K}\mathcal{S}(f_{i},s_{j})\right)\geq 1}\Bigg\}}{m}$ (1)
+Not only does it evaluate if a model can find support sentences, but it also evaluates if a model can control information redundancy from these sentences. Figure 5 in Appendix C visualizes SSR.
+Bias Type
+Model Avg Selection Attrition Performance Detection Reporting
+n = 157 n = 46 n = 54 n = 61 n = 14
+OpenAI-v3 22.7 40.1 6.3 26.7 27.4 13.1
+GritLM-7B 18.9 35.8 7.0 15.2 26.8 9.5
+GPT-4o 47.5 60.5 42.4 41.5 50.1 43.0
+Sonnet-3.5 39.2 55.4 30.7 37.0 37.2 35.5
+Llama-3.1-70B 45.6 61.2 34.2 45.2 50.1 37.4
+Llama-3-8B 22.7 49.4 14.5 22.1 20.8 6.5
+Llama-3-8B 40.8 69.0 24.8 47.3 48.6 14.3
+Fine-tuned
+Table 3: Support Sentence Retrieval (SSR).
+Evaluated on Cochrane Test (SSR).
+Metric is Aspect Recall Ratio @ Optimal.
+Llama-3-8B fine-tuned on Cochrane Train (SSR).
+Fine-tuning details are included in Appendix E.
+Bias Type
+Model Avg Selection Attrition Performance Detection Reporting
+n = 130 n = 98 n = 87 n = 82 n = 80
+GPT-4o 47.2 58.5 60.2 48.3 42.7 26.3
+Sonnet-3.5 59.9 73.1 73.5 50.6 51.2 51.3
+Llama-3.1-70B 53.2 66.2 62.2 44.8 46.3 46.3
+Llama-3-8B 26.5 26.9 34.7 24.1 22.0 25.0
+Llama-3-8B 29.6 40.8 28.6 24.1 19.5 35.0
+Fine-tuned
+Table 4: Support Judgment Selection (SJS). Evaluated on Cochrane Test (SJS). Metric is Accuracy. Llama-3-8B fine-tuned on Cochrane Train (SJS). Fine-tuning details are included in Appendix E.
+Bias Type
+Model Avg Selection Attrition Performance Detection Reporting Deviation
+n = 933 n = 629 n = 309 n = 645 n = 594 n = 331
+GPT-4o 42.1 50.5 36.5 54.7 43.7 34.4 32.7
+Sonnet-3.5 41.9 52.8 37.1 50.6 43.3 33.6 34.2
+Llama-3.1-70B 38.8 48.1 31.7 48.0 44.4 31.6 29.0
+Llama-3-8B 30.1 36.4 32.4 39.1 37.2 19.8 15.4
+Llama-3-8B Fine-tuned 36.3 49.5 33.3 42.4 40.1 27.0 25.7
+LR Trained N/A 26.0 21.7 27.9 28.6 21.8 N/A
+SVM Trained N/A 30.7 24.2 23.3 25.6 23.5 N/A
+Table 5: Main Task, Risk-of-Bias Determination.
+Evaluated on Cochrane Test (Main) + Non-Cochrane Test (Main).
+Metric is Macro-F1.
+Llama-3-8B, logistic regression, and SVM are fine-tuned/trained on Cochrane Train (Main).
+Fine-tuning details are included in Appendix E.
+4.5 Support Judgment Selection (SJS)
+Simply retrieving support sentences is insufficient for a risk-of-bias determination.
+A model needs the ability to interpret and synthesize information from these retrieved support sentences in order to generate a good support judgment that will lead to the correct determination of risk-of-bias.
+Following the convention of using multiple-choice questions as a proxy for generative tasks Rein et al.
+(2023), we propose the support judgment selection subtask, an MCQ task, where the model selects the correct support judgment from a mix of three synthetically generated support judgments, and three human-written options derived from other papers’ support judgments for the same exact bias.
+These three human-written options sometimes proved to be surprisingly hard distractors because their lack of direct paper quotes made them into generic, non-paper-specific statements.
+A deep logical understanding of the paper’s content rather than superficial semantic matching is required to look past them and find the correct one.
+We prompt GPT-4 to generate the three synthetic options by imitating support judgments from other papers concerning the same exact bias name.
 These options are tailored to be paper-specific while maintaining the underlying reasoning.
-See Appendix D for details on construction of these synthetic judgments.
-In order to ensure that all incorrect options were actually incorrect, we performed filtering in two stages.
-GPT-4 was used to perform preliminary filtering (the prompt is shown in Figure 13 of Appendix D).
-In the second stage, we manually reviewed all of the remaining options, and removed any that were not incorrect.
-Figure 4 shows an example of a multiple-choice question with one correct answer and three synthetically generated answers.
-Both options C and D refer to the same information from the paper, in this case the operational challenges affecting the trial’s methodology, but describe different reasoning given this information.
-This demonstrates that retrieving the correct information from the paper is not sufficient for solving this task.
-Figure 4: Risk of Bias Support Judgment Selection.
-In Task 3, the model is shown a paper, and asked which of 7 support judgments provides the best explanation of the paper’s risk of bias.
-(Only four options are shown here for illustration purposes.)
-3.5 Task 4: Risk Level Determination
-The final task directly evaluates a model’s ability to assess the risk of bias in biomedical research papers.
-It requires the model to categorize the risk level of a specific bias as high, low, or unclear.
-The input to the model includes the whole paper, the PICO of the study, the objective of the meta-analysis, and the definition of the bias, as illustrated in Figure 5.
-The review authors use bias with definitions from Cochrane Handbook 5.1 Higgins and Green [2011], Cochrane Effective Practice and Organisation of Care Practice and of Care [EPOC], A Cochrane Risk of Bias Assessment Tool Sterne et al.
-[2014], or specific biases that are defined within the meta-analysis.
-For the last category, we extract the bias definition from the specific meta-analysis.
-(Bias definitions are also extracted for Tasks 2 and 3.)
-This assessment requires synthesizing various elements identified in the previous tasks, including the accurate retrieval and interpretation of relevant sentences and the ability to critically evaluate the sufficiency and reliability of the evidence presented regarding each bias.
-Figure 5: Risk Level Determination. In Task 4, the goal is to provide an assessment of the paper’s risk of specific biases.
+Empirically, these six distractor options are often misleading.
+Table 4 shows the best LLM only achieves 60% accuracy.
+In order to ensure that all incorrect options were actually incorrect, all 906 datapoints were checked by human, and 465 datapoints were kept.
+See Figure 6 in Appendix C to visualize SJS.
 5 Experiments
-4.1 Experimental Models and Procedures
-We evaluate GPT-4o OpenAI [2024], Claude3-Opus Anthropic [2024a], Claude3.5-Sonnet Anthropic [2024b], and Gemini-1.5-Pro Google [2024].
-Due to the long-context nature of RoBBR’s tasks, with many datapoints exceeding 16k tokens, many open-sourced LLMs AI [2024], Poli et al.
-[2023], Li et al.
-[2023] are considered but none of which is equipped to evaluate RoBBR.
-We use chain-of-thought as prompting strategy to stabilize the model performance. We optimize our prompts on a separate development set. See Appendix E for details on prompt optimization.
+5.1 Evaluated Models
+We evaluate various LLMs, GPT-4o-2024-05-13 OpenAI (2024a), Sonnet3.5-20240620 Anthropic (2024), and Llama-3.1-70B Dubey et al.
+(2024).
+We evaluate and fine-tune light-weight Llama-3-8B Dubey et al.
+(2024).
+We also evaluate two recent embedding models GritLM 7B Muennighoff et al.
+(2024) and OpenAI-embedding-v3-large OpenAI (2024).
+Note the GPT-4o model is different from the GPT-4-0125 model used in our annotation pipeline.
+For the main task, we further evaluate three traditional ML models: RobotReviewer Marshall et al. (2016), Support Vector Machine (SVM) and Logistic Regression (LR) Dias et al. (2025).
+For all generative language models, we use chain-of-thought as a prompting strategy to stabilize the model performance.
+We develop our prompts on train sets.
+For reproducibility, we use a common generation setting of temperature 0, top-p 1, frequency penalty 0, and presence penalty 0.
 5.2 Evaluation Metrics
-For Task 1, study inclusion/exclusion, we report macro-F1, accuracy, and weighted accuracy.
-For Task 2, Risk of Bias Sentence Retrieval, we report aspect-level recall metrics, which measure the proportion of the bias aspects that are covered by the retrieved sentences.
-Specifically, Bias Recall (BR) @ 3 measures the proportion aspects covered when three sentences are retrieved.
-BR @ Optimal measures recall when the “optimal" number of sentences are retrieved.
-The optimal number of sentences is defined as the fewest which are sufficient to cover all bias aspects.
-For Task 3, Risk of Bias Support Judgment Selection, we report accuracy.
-For Task 4, Risk Level Determination, we report accuracy and weighted accuracy.
-4.3 Categorization
-For Tasks 2, 3, and 4, we consider 8 categories of biases: selection bias, attrition bias, performance bias, detection bias, reporting bias, analysis bias, special bias (the biases that are meta-analysis specific), and other bias (i.e.
-biases that do not fall into any of the other categories).
-In order to reduce evaluation costs, we randomly sampled 50 task instances for some categories.
-It is important to note that systematic reviews vary in which biases they report.
-When the authors judge that a bias is not important for the objective of a review, they may decide not to evaluate a specific bias.
-As a result, not all categories have the same number of examples, and some categories have fewer than 50 examples.
-See Table 10 for detailed statistics.
-5 Results
-Table 2: Results for Task 1: Study Inclusion/Exclusion
-F1 Accuracy Weighted Accuracy
-Claude3-Opus 75.5 77.1 74.9
-Claude3-Sonnet 80.6 81.1 80.8
-Gemini-1.5-Pro 72.2 72.7 73.0
-GPT-4o 75.5 75.6 77.6
-[…]
-Bias Type
-Model Full Selection Attrition Performance Detection Reporting Analysis Special Other
-n = 337 n = 117 n = 57 n= 50 n = 50 n = 35 n = 13 n = 7 n = 23
-Bias Recall@Optimal
-Claude3-Opus 39.0 55.4 22.0 57.3 46.4 11.0 47.9 13.9 4.4
-Gemini-1.5-Pro 41.1 57.2 28.7 49.1 48.8 23.3 54.2 13.9 3.3
-GPT-4o 44.8 60.3 28.6 64.6 48.7 32.9 28.7 28.0 2.3
-Bias Recall@3
-Claude3-Opus 52.3 67.9 40.2 63.1 55.9 25.5 55.9 64.6 19.7
-Gemini-1.5-Pro 56.1 69.7 44.1 73.7 65.3 35.6 56.3 35.5 6.7
-GPT-4o 61.8 72.2 47.9 82.0 78.5 43.2 67.9 50.4 10.3
-5.1 Performance on the Four Tasks
-In Task 1 (Study Inclusion/Exclusion), shown in Table 2, Claude3-Opus and GPT-4o perform best in terms of F1 score. Claude3-Opus has higher accuracy and GPT-4o higher weighted accuracy.
-In Task 2 (Sentence Retrieval Evaluation Results) 3, GPT-4o has stronger results than both Gemini-1.5 and Claude3-Opus for almost all bias types.
-However, in Task 3 (Support Judgment Selection Evaluation Results) 4, Claude3-Opus has the strongest results.
-Tasks 2 and 3 are critical for evaluating the different capabilities of the LLMs.
-Task 2 is a long-context task, requiring retrieval of a non-redundant set of sentences from a 200-sentence paper.
-Task 3 requires complex reasoning to evaluate the retrieved sentences and identify the implications for risk of bias.
-Our results demonstrate that GPT-4o performs better in long-context retrieval tasks, while Claude3-Opus exhibits stronger reasoning abilities.
-Figure 4 illustrates Claude3-Opus’s stronger reasoning skills.
-Options C and D present different reasoning based on the same information from the paper.
-Notably, GPT-4o is misled by a deceptive option, whereas Claude3-Opus correctly handles the multiple-choice question.
-Task 4 (Risk Level Determination Evaluation Results) 5 provides an overall assessment of each model’s ability to determine the risk level, integrating all information from the previous tasks.
-The three models show comparable performance in this task.
-[…]
-Bias Type
-Model Full Selection Attrition Performance Detection Reporting Analysis Special
-n = 310 n = 86 n = 60 n = 52 n = 56 n = 50 n = 8 n = 6
-Accuracy
-Claude3-Opus 51.2 62.6 56.7 48.0 46.5 44.2 50.0 0.0
-Claude3.5-Sonnet 60.2 69.8 79.9 55.8 44.8 51.8 62.5 0.0
-Gemini-1.5-Pro 49.3 57.1 59.9 46.4 42.8 40.2 49.3 16.6
-GPT-4o 45.8 52.3 63.1 42.5 37.7 28.0 74.5 0.0
-[…]
-Bias Type
-Model Full Selection Attrition Performance Detection Reporting Analysis Special Other
-n = 382 n = 93 n = 50 n = 50 n = 50 n = 50 n = 18 n = 26 n = 50
-F1
-Claude3-Opus 47.7 54.3 34.1 46.7 37.4 23.0 69.9 27.9 27.9
-Claude3.5-Sonnet 49.7 62.7 23.9 49.9 46.7 30.1 48.3 28.7 33.8
-Gemini-1.5-Pro 43.1 53.9 26.0 43.7 46.8 30.4 44.6 20.1 22.8
-GPT-4o 51.9 59.0 36.5 48.9 36.0 27.7 53.8 34.7 29.8
-Accuracy
-Claude3-Opus 56.8 64.6 58.1 48.2 41.7 40.0 88.9 53.8 72.2
-Claude3.5-Sonnet 62.3 64.4 52.0 52.1 55.9 64.3 83.0 69.5 72.1
-Gemini-1.5-Pro 49.3 62.5 42.4 46.5 48.1 56.0 77.5 42.4 23.7
-GPT-4o 62.6 71.0 68.3 54.4 40.1 57.6 88.7 65.4 62.1
-Weighted Accuracy
-Claude3-Opus 49.5 55.3 36.2 48.0 46.5 30.4 69.3 28.2 32.4
-Claude3.5-Sonnet 50.3 63.1 24.1 51.6 48.2 32.1 46.8 36.1 36.1
-Gemini-1.5-Pro 47.9 56.3 34.0 44.4 54.2 36.4 43.6 22.2 32.5
-GPT-4o 53.6 60.4 36.3 51.3 40.7 30.0 50.0 64.3 32.6
-5.2 Performance across Bias Types
-The three models exhibit significant variations in performance across different biases due to the varying challenges associated with each bias type.
-All models have stronger performance at assessing selection bias, as biomedical papers typically detail their randomization processes explicitly.
-In contrast, assessing reporting bias requires a comprehensive understanding of the entire paper to determine whether all outcomes specified in the study protocol are reported.
-Evaluating attrition bias involves extracting data on participant dropout during the trial and assessing how this dropout impacts the study’s statistical estimates.
-The most challenging bias type to assess is "other bias," where the model must grasp all potential biases, such as those stemming from funding sources or bias due to lack of power.
-Successfully identifying them requires a thorough understanding of all possible biases a study might have, thereby necessitating the treatment of each bias type as a distinct task.
-6 Related Work
-Scientific and Biomedical Information Retrieval has been the subject of a significant body of work in recent years Jin et al.
-[2023], MacAvaney et al.
-[2020], Wang et al.
-[2023].
-A subset of this work has been focused on benchmarks Ben Abacha and Demner-Fushman [2019], Krithara et al.
-[2023], Voorhees et al.
-[2021], Cohen et al.
-[2017].
-Study Screening Many works have focused on the task of screening studies for inclusion in systematic reviews Cohen et al.
-[2006], Khabsa et al.
-[2016], Kontonatsios et al.
-[2017], El-Gayar et al.
-[2015], with more recent work often focusing on LLMs Wilkins [2023], Tran et al.
-, Guo et al.
-[2024], Na et al.
-[2024], Cai et al.
-[2023], Ye et al.
-[2024], Wang et al.
-[2024].
-Huotala et al.
-[2024] conduct an evaluation of LLM paper screening, finding that few-shot chain-of-thought prediction performs comparably to human annotators.
-Robinson et al.
-[2023] leverage an instruction-tuned Guanaco-7B model on a large dataset of systematic reviews to exceed the screening performance of ChatGPT.
-Cohen et al.
-[2006] produce a dataset of screening judgments for 15 systematic drug reviews.
-Study Risk of Bias Risk of bias identification requires a thorough reading of an entire paper.
-Marshall et al.
-[2015, 2016] train an SVM model with documents annotated for risk of bias, creating RobotReviewer system to predict risk of bias levels.
-Wang et al.
-[2022b] utilized a BERT-based pipeline for risk of bias prediction.
-However, several other works find lower agreement rates between LLMs and human annotators Lai et al.
-[2024], Hasan et al.
-[2024], Barsby et al.
-[2024].
-Other LLM Extraction Efforts Recent years have seen LLMs benchmarked in extracting "ground truth" from systematic reviews, particularly in the context of data extraction Sun et al.
-[2024], Gartlehner et al.
-[2023], multi-document summaries Wang et al.
-[2022a], Wallace et al.
-[2020] and relevance prediction Al-Hussaini et al.
-[2022].
-Another use case has been the extraction of PICOs (Participants, Interventions, Comparisons, Outcomes) from studies.
-While earlier methods achieved strong results Zafar et al.
-[2023], Hu et al.
-[2023], Wallace et al.
-[2016], Dhrangadhariya and Müller [2023], Jin and Szolovits [2020], Kang et al.
-[2019], a body of recent work has focused on extracting PICOs with LLMs.
-Fine-tuning Wadhwa et al.
-[2023] and prompt tuning Tang et al.
-[2024] have been used to enhance extraction.
-Evaluation of Zero-shot LLM Extraction Kartchner et al.
-[2023] show that ChatGPT performed inconsistently on various meta-analysis extraction tasks.
-For instance, while it correctly extracted cancer type roughly 90% of the time, treatment type accuracy fell below 25%.
-This inconsistency is a considerable motive to more comprehensively understand LLM reliability on various meta-analysis extraction tasks.
-Limitation Tasks in Scientific Documents Early efforts in assessing papers’ limitations focused on performing summary extraction and evaluating the result using human evaluation Cohan et al.
-[2018, 2022], Liu and Shah [2023].
-For understanding scientific documents, identifying and retrieving a paper’s limitation is essential to understand the utility and viability of claims and evidences presented in the paper.
-Faizullah et al.
-[2024] evaluate fine-tuned embedding models and LLMs’ ability to generate limitation using the extracted limitation from the original paper.
-Marshall et al.
-[2016] use direct quotes in risk of bias judgment to obtain annotations for review authors’ judgments, where exact matching texts are labeled as relevant to a risk of bias.
-Suster et al.
-[2023] present a quality assessment task to assess a biomedical paper using the Grading of Recommendation, Assessment, Development, and Evaluation with justifications.
+Some bias names may fall into multiple categories, causing the total datapoint count across categories to be slightly higher than the test set size.
+Model performance is reported per category and as an average across all categories.
+The SSR subtask uses Aspect Recall Ratio @ Optimal defined in Equation 1 as its metric.
+It measures the percentage of aspects covered by the retrieved sentences.
+Unlike traditional metrics (e.g., recall, precision), it penalizes redundancy by rewarding models that retrieve the minimum number of sentences (typically 1-2) needed to cover all aspects for a given data point.
+For instance, if one sentence covers multiple aspects, retrieving additional sentences that cover already addressed aspects is discouraged.
+6 Benchmarking Results and Analyses
+6.1 LLMs Exhibit Distinct Strengths in Retrieval and Reasoning
+While GPT-4o and Sonnet-3.5, with 42.1% and 41.9% average Macro-F1, are comparable in the main task Table 5, a closer inspection of their performances on the two subtask Tables 4, 4 reveals interesting differences.
+In support sentence retrieval (SSR), GPT-4o has 47.5% Aspect Recall Ratio, while Sonnet-3.5 only has 39.2%.
+This shows that Sonnet-3.5 is worse at retrieving the best support sentences that are the bases of human expert-level support judgments.
+However, in support judgment selection (SJS), Sonnet-3.5 outperforms other GPT-4o with 59.9% over 47.2% accuracy.
+This shows that Sonnet-3.5 can select better support judgments, which involves reasoning and deliberating which option best aligns with the paper’s contents and the bias and risk-level definitions.
+In other words, GPT-4o’s advantage over Sonnet-3.5 is information retrieval, while Sonnet-3.5’s advantage is information synthesis with reasoning.
+We provide an example in Appendix G comparing reasoning traces of the two models to clearly illustrate Sonnet-3.5’s advantage in reasoning.
+6.2 Positive Correlation Between Main Task and Subtasks
+We investigate the relationship between the main task and our subtasks.
+For an LLM (either GPT-4o or Sonnet 3.5), we categorize datapoints in the SSR subtask into two classes: datapoints whose risk-of-bias level is correctly predicted by the LLM and those incorrectly predicted.
+For each of the two classes, we calculate its average Aspect Recall Ratio @ Optimal.
+See the left part in Figure 4, we observe that the Aspect Recall Ratio is significantly higher in the correct class than the incorrect class, for both Sonnet-3.5 and GPT-4o, suggesting LLMs’ performances on SSR are positively correlated with their performances on the main task.
+Similarly, datapoints in the SJS subtask are divided into two classes.
+From the right part in Figure 4, we see that, for both Sonnet-3.5 and GPT-4o, their performances on SJS are positively correlated with their performances on the main task.
+Figure 4: There is a positive correlation between the main task and SSR & SJS subtask performance.
+6.3 Retrieval and Reasoning Ablation
+Retrieval Reasoning Main Task Result
+Sonnet-3.5 GPT-4o 31.5 $\pm$ 5.0
+GPT-4o GPT-4o 40.3 $\pm$ 6.9
+Ground Truth GPT-4o 46.1 $\pm$ 7.5
+Ground Truth Sonnet-3.5 49.4 $\pm$ 8.0
+Table 6: Two stage pipeline that first retrieves and then reason based on the retrieved results (full paper is not shown).
+Main Task, Risk-of-Bias Determination.
+Evaluated on the 313 datapoints in Cochrane Test (SSR), all of which have risk-of-bias labels.
+Metric is Macro F1 score (%) averaged across bias categories.
+The error bar represents the 95% confidence interval.
+To investigate how the retrieval and reasoning abilities of LLMs affect their performance on the main task, we employ a two-stage retrieval and reasoning pipeline.
+In this setup, one LLM processes sentences retrieved by another LLM, rather than accessing the original paper directly.
+The specific prompt used is detailed in Appendix I.3.
+This experiment was conducted on Cochrane Test (SSR), which comprises 313 datapoints, each associated with a support judgment and a risk-of-bias label.
+We pair different models’ retrieved sentences with different reasoning models to evaluate performance on the main task Risk-of-Bias Determination.
+Ground truth retrieved sentences are provided by our annotation pipeline.
+As shown in Table 6, when results from different retrieval methods are paired with the same reasoning model (GPT-4o), ground-truth retrieval achieves significantly higher macro F1 score (46.1%) over Sonnet-3.5’s retrieval (31.5%), confirming that retrieval quality impacts main task performance.
+When the retrieval quality is held constant (using ground-truth retrieval) and the reasoning model varied, there is some inconclusive evidence that suggests Sonnet-3.5 slightly outperforms GPT-4o.
+These ablations demonstrate that both retrieval and reasoning abilities are relevant for the Risk-of-Bias Determination task.
+6.4 Retrieval Marginally Improves LLM on Non-Cochrane Test Set
+So far we have focused on the Cochrane test set, which includes support judgments, unlike the Non-Cochrane test set.
+To investigate whether more accurate support sentences could enhance LLM performance on the Non-Cochrane test set (2489 datapoints), we used the o4-mini reasoning model OpenAI (2024b) to find such sentences within each paper, guided by the datapoint’s actual risk level (see Appendix I.4 for the prompt).
+In this proof-of-concept setup, the LLM received the full paper and o4-mini-identified support sentences as highlights, which we compared against baselines with full paper but no highlights.
+Table 7 reveals that both GPT-4o and Sonnet-3.5 showed slight performance improvements on the Non-Cochrane Test (Main) over their baselines.
+The limited gain is partly due to inaccuracies in retrieved sentences, as recovering support sentences without expert support judgments is challenging.
+Setting Main Task
+GPT-4o Full Paper 42.5 $\pm$ 2.2
+GPT-4o Full Paper + Highlight Retrieval 44.1 $\pm$ 2.2
+Sonnet-3.5 Full Paper 40.3 $\pm$ 2.2
+Sonnet-3.5 Full Paper + Highlight Retrieval 44.5 $\pm$ 2.2
+Table 7: Main Task, Risk-of-Bias Determination.
+Evaluated on Non-Cochrane Test (Main).
+Metric is Macro F1 score (%) averaged across bias categories.
+The error bar represents the 95% confidence interval.
+6.5 Challenges in Support Sentence Retrieval
+Our findings establish the retrieval task SSR as a crucial intermediate milestone for improving LLM on the main task.
+Nevertheless, Table 4 indicates that in zero-shot setting, current models struggle significantly: modern embedding models such as OpenAI-v3 and GritLM-7B achieve low recall ($\leq$22.7%), while commercial LLMs also underperform (recall $\leq$47.5%).
+Although fine-tuning LLama3-8B boosts its SSR recall from 22.7% to 40.8%, its retrieval quality remains unsatisfactory.
+A significant limitation of embedding models is their lack of context awareness.
+Even recent contextual-aware models Morris and Rush (2024) cannot handle the long-range interactions across different sentences.
+In Appendix F, we examine a type of globally-connected datapoints in SSR potentially necessitates a more global understanding of the entire context.
+We leave the exploration of contextual embedding models and more accurate long-context LLMs for solving SSR as a future research direction.
+6.6 Limitations of Traditional ML Models on RoBBR
+Biases
+Model Avg Allocate Blind Blind RanSeq
+Conceal Outc Part Gen
+n=32 n=19 n=18 n=30
+RobotReviewer 56.7 $\pm$ 8.4 75.0 39.1 43.8 68.9
+LR 53.1 $\pm$ 9.7 71.9 50.4 51.8 38.4
+SVM 44.8 $\pm$ 8.6 45.9 55.2 41.9 36.0
+GPT-4o 65.6 $\pm$ 8.5 83.6 59.1 41.9 77.8
+Sonnet-3.5 67.5 $\pm$ 8.4 77.0 82.5 41.9 68.8
+Table 8: Main Task, Risk-of-Bias Determination.
+Evaluated on a subset from the Cochrane Test (Main), comprising only the four RobotReviewer-assessable bias types: allocation concealment, blinding of outcome, blinding of participants, and random sequence generation.
+Only two judgment categories (low, and high/unclear) per RobotReviewer’s specification.
+Metric is Macro-F1.
+The error bar represents the 95% confidence interval.
+This study compares previous ML models with LLMs on four specific biases, the only ones assessable by RobotReviewer Marshall et al.
+(2017) (an SVM/CNN ensemble) due to its inaccessible training code.
+We also trained logistic regression and SVM models on our Cochrane Train Set (Main) Dias et al.
+(2025).
+Table 8 shows inconclusive results : RobotReviewer, SVM, and LR perform marginally worse, if at all, than GPT-4o and Sonnet-3.5.
+The four biases RobotReviewer assesses are considered straightforward because superficial keywords often directly indicate bias risk (e.g., "opaque envelope" for low allocation concealment bias; "randomness" or "random number generator" for low random sequence generation bias), suggesting significant keyword reliance.
+However, biases like selective reporting or deviation bias may necessitate a more holistic consideration of the entire biomedical paper.
+Traditional ML models by definition are hard to generalize to new bias names and definitions.
+RoB1/RoB2 guidelines list hundreds of bias names, some with definitions re-interpreted in different reviews.
+we treat these as paper-specific biases, manually incorporating the paper’s re-interpreted definition during LLM evaluation on RoBBR.
+Thus, RoBBR primarily evaluates models in zero-shot settings or fine-tunes models like Llama3-8B for generalization to unseen bias definitions via semantic understanding.
+Furthermore, most bert-based models’ 512-token context limit is ill-suited for RoBBR’s biomedical studies (average 8k tokens), unlike modern LLMs with longer context windows.
 7 Conclusion
-We have presented RoBBR, a benchmark for assessing the ability to judge risk of bias in biomedical papers.
-The benchmark is based on gold-standard risk of bias assessments from Cochrane, and uses a human-validated annotation pipeline for scalably constructing task examples.
-The four benchmark tasks evaluate different skills required for these assessments, including retrieval and reasoning.
-Performance of state-of-the-art language models on the benchmark tasks reveals that current models fall significantly short of expert-level performance.
-By providing a standardized tool for measuring risk of bias judgments, the RoBBR benchmark can help guide the development of AI systems that aim to automatically assess study quality and synthesize scientific evidence.
-As systems that review the scientific literature are developed, it is important that they can reliably judge the strength of evidence in order to draw sound conclusions.
-ML systems which perform well on this benchmark could potentially reduce the time required for risk of bias assessments in systematic reviews, which currently averages nearly six hours per study Crocker et al.
-[2023].
-We hope the RoBBR benchmark will spur further research into developing trustworthy systems for this task.
-8 Limitations
-The RoBBR benchmark has several limitations.
-The source of all bias judgments in the benchmark is Cochrane meta-analyses.
-While these are considered very high quality meta-analyses among biomedical researchers, the bias judgments that they contain are likely to have occasional errors.
-In addition, there can be disagreements in judgment among different meta-analysis authors.
-The Cochrane risk of bias framework does not necessarily identify all types of bias in research studies.
-The framework uses a predefined set of bias categories, which may not be appropriate for all studies.
-The benchmark only includes interventional studies, and therefore excludes observational studies.
-Assessment of risk of bias in observational studies requires a different procedure than for interventional studies.
-Our benchmark does not measure performance for observational studies.
-Only a limited number of prompting strategies were evaluated for the LLMs. It is likely that improved prompting and task presentation would improve performance on the tasks.
-Appendix: Table of Contents
-  1. A.
-    DatasetA
-    1. 1.
-      Dataset License and Code LicenseA.1
-    2. 2.
-      Dataset Hosting, Accessibility and MaintenanceA.2
-    3. 3.
-      Dataset StatisticsA.3
-    4. 4.
-      Dataset Collection and Processing A.4
-    5. 5.
-      RoBBR Structure A.5
-  2. B.
-    Annotation GuidelineB
-  3. C.
-    Task 2 Bias Retrieval Prompt and OptimizationC
-    • Aspect Decomposition Prompt C.1
-    • Aspect Filtering Prompt C.2
-    • Aspect-to-Sentence Mapping Prompt and Optimization C.3
-    • A Motivating Example C.4
-  4. D.
-    Task 3 Synthetic Option Generation Procedure and PromptD
-  5. E.
-    Experiment DetailsE
-    • Prompt Optimization E.1
-    • Prompts for Evaluation E.2
-      • Prompt to Evaluate Study Inclusion/Exclusion (Task 1) E.2.1
-      • Prompt to Evaluate Bias Retrieval (Task 2) E.2.2
-      • Prompt to Evaluate Support Judgment Selection (Task 3) E.2.3
-      • Prompt to Evaluate Risk Level Determination (Task 4) E.2.4
-    • Results with Standard Error E.3
-  6. F.
-    Full Example F
-    • Search Protocol F.1
-    • Objective of the Meta-Analysis F.2
-    • Full Paper F.3
-    • Bias Name and Definition F.4
-    • PICO F.5
+We present RoBBR, a benchmark for measuring models’ ability to assess risk-of-bias in biomedical studies.
+RoBBR includes two novel subtasks that evaluate a model’s retrieval and reasoning abilities.
+The support sentence retrieval subtask is created by our fully automatic and human-validated annotation pipeline for aligning support judgments to sentences in biomedical studies.
+Our analysis reveals the importance of retrieval and reasoning abilities and demonstrate their impact on the a model’s ability to assess risk-of-bias.
+Limitations
+While in this work all systematic reviews and biomedical studies are open-sourced (Creative Commons License or Public Domain), many other systematic reviews have licenses which restrict distributing their contents.
+Therefore, we share the entire codebase for creating the RoBBR benchmark and encourage researchers with access to non-open-source contents to create other versions of RoBBR.
+In our evaluation of LLMs, while we have developed our prompts using RoBBR’s train sets and investigated popular techniques such as chain-of-though reasoning etc, we acknowledge the possibility that better prompting techniques could lead to slight model improvements on RoBBR.
+Ethics Statement
+All data in RoBBR come from open-sourced systematic reviews and biomedical studies covering the domain of evidence-based biomedicine.
+The authors make sure no personal information is included in RoBBR by manual inspection of all datapoints.
+The RoBBR benchmark, which promotes the automation of risk-of-bias determination, might introduce over-reliance on commercial AI systems to evaluate a biomedical study’s risk-of-bias.
+The benchmark only includes systematic reviews and biomedical studies in English.
+Systems trained and evaluated on RoBBR might disadvantage the risk-of-bias determinations for non-English biomedical studies.
 Appendix A Dataset
 A.1 Dataset License and Code License
-The RoBBR dataset is made available under CC-BY-NC. A copy of the full license can be found at https://github.com/RoBBR-Benchmark/RoBBR/blob/main/LICENSE.md.
+The RoBBR dataset is made available under CC-BY-NC. A copy of the full license can be found at RoBBR License.
 The code used in this paper is released under the MIT License.
 The MIT License is a permissive open-source license that allows for the free use, modification, and distribution of the code, as long as the original license is included with any derivative work.
-A copy of the full license can be found at https://github.com/RoBBR-Benchmark/RoBBR/blob/main/LICENSE.md.
+A copy of the full license can be found at RoBBR License.
 A.2 Dataset Hosting, Accessibility and Maintenance
-The RoBBR dataset with its meta-data is released and can be accessed freely at (https://github.com/RoBBR-Benchmark/RoBBR/blob/main/LICENSE.md).
+The RoBBR dataset with its meta-data is released and can be accessed freely at RoBBR License.
 We commit to regularly maintain the dataset and codebase by incorporating user feedback.
@@
 A.3 Dataset Statistics
-See table 6 7 8 9 for detailed test set statistics of the four tasks.
-Table 6: Test set statistics for Task 1 (Study Inclusion/Exclusion)
-# of Tokens in Query Correct Option Count
-Dataset Points min avg max Included Excluded
-Test Set 309 2685 9380 28464 128 181
-Table 7: Test set statistics for Task 2 (Risk of Bias: Sentence Retrieval)
-# of Tokens in the Query Sentences Covered Aspects Optimal # of sentences
-Dataset Points min avg max min avg max min avg max min avg max
-Test Set 513 4104 11220 21494 87 220.1 400 1 1.9 10 1 1.4 9
-Table 8: Test set statistics Task 3 (Risk of Bias: Support Judgement Selection)
-# of Tokens in the Query
-Dataset Points min avg max
-Test Set 479 3515 10583 21187
-Table 9: Test set statistics for Task 4 (Risk of Bias: Risk Level Determination)
-# of Tokens in Query Correct Option Count
-Dataset Points min avg max Low Unclear High
-Test Set 1052 3256 9763 19834 676 213 163
-Table 10: Test set category statistics for Task 2, 3, and 4
-Bias Type
-Task Points Selection Performance Detection Reporting Attrition Other Analysis Paper specific
-Task 2 513 202 81 82 35 95 23 13 7
-Task 3 479 130 87 82 80 98 0 8 6
-Task 4 1052 332 170 168 119 148 102 18 26
+See Tables 9, 10, 11 for detailed statistics of the RoBBR tasks.
+Statistic min avg max
+Tokens per report 2,908 9,215 19,080
+Sentences 87 218.6 400
+Covered Aspects 1 1.9 7
+Optimal # of sentences 1 1.4 5
+Table 9: Statistics for the Cochrane Test Set SSR (313 data points).
+Bias Type Cochrane Test SSR Cochrane Test SJS Cochr. + Non-Cochr. Test Main
+Selection 157 130 933
+Attrition 46 98 629
+Performance 54 87 309
+Detection 61 82 645
+Reporting 14 80 594
+Deviation 0 0 331
+Table 10: Test set category statistics
+Statistic Cochrane Train Non-Cochr. Test Cochrane Test
+Points 774 2 489 906
+Tokens per report:
+min 1,991 1,488 2,274
+avg 9,907 7,791 8,946
+max 18,894 26,154 19,080
+Label distribution:
+Low 387 1,465 562
+Unclear 275 574 195
+High 112 450 149
+Table 11: Test-set statistics for Risk-of-Bias Determination. Unclear/Some concerns for Non-Cochrane test
 A.4 Dataset Collection and Processing
-[…]
+We use BioC Comeau et al.
+(2019) API to download meta-reviews and papers in PMC database.
+We manually download the rest of the papers.
+We use GROBID Lopez (2008–2024) to parse papers from PDF to XML format.
+We use Stanza Qi et al.
+(2020) to split paragraphs (including table and figure captions) into sentences.
 A.5 RoBBR Structure
-  • task1_SIE_test.json and task1_SIE_dev.json: Test and dev set of task 1 Study Inclusion/Exclusion
-    • paper_doi: The DOI of the paper.
-    • objective: The meta-analysis objective.
-    • search_protocol: search protocol information of the meta-analysis.
-    • full_paper: The full paper content.
-    • label: One of [included, excluded], showing whether the paper is included or excluded in the meta-analysis.
-  • task2_ROBSR_test.json and task2_ROBSR_dev.json: Test and dev set of task 2 Risk of Bias Sentence Retrieval
+  • SSR_test.json and SSR_dev.json: Test and dev set of Support Sentence Retrieval
     • paper_doi: The DOI of the paper.
@@
     • sentence_index2aspects: a mapping (i.e. dictionary) between sentence index and all aspect ids that this sentence is the source of information of.
-    • bias_retrieval_at_optimal_evaluation: This is a dictionary that contains the necessary information for evaluating the model’s performance on the task Bias Retrieval @Optimal.
+    • bias_retrieval_at_optimal_evaluation: This is a dictionary that contains the necessary information for evaluating the model’s performance on the task Support Sentence Retrieval @ Optimal.
         optimal: A positive integer, which is the smallest number of sentences needed to cover the largest number of aspects.
@@
         covered_aspects: the list of aspects that are covered. In this case, the list of aspects covered is list of all aspects.
-    • bias_retrieval_at_3_evaluation: This is a dictionary that contains the necessary information for evaluating your model’s performance on the task Bias Retrieval @3.
-        one_selection_of_sentences: a list of 3 sentence indices.
-The list of sentences cover the largest number of aspects that can be covered under the restriction of 3 sentences.
-Note, there are potentially other lists of sentences that has size 3 and cover the same number of aspects.
-        covered_aspects: the list of aspects that are covered.
-In this case, this list of aspects may not be all the aspects.
-Since in the paper, we calculate aspect recall by diving the number of aspects covered by model’s retrieved sentences against the total number of aspects, for Bias Retrieval @3, the maximum possible performance is not 100%.
-  • task3_SJS_test.json and task3_SJS_dev.json: Test and dev set of task 3 Support Judgment Selection
+  • SJS_test.json and SJS_dev.json: Test and dev set of Support Judgment Selection
     • paper_doi: The DOI of the paper.
@@
     • label: The index of the correct option.
-  • task4_RLD_test.json and task4_RLD_dev.json: Test and dev set of task 4 Risk Level Determination
+  • Main_task_test.json and Main_task_dev.json: Test and dev set of the main task
     • paper_doi: The DOI of the paper.
@@
   • Ultimately, you and your teammate should collaboratively arrive at a consensus for each of the 50 tasks. Write the collective answers in the answer file provided.
-Appendix C Task 2 Bias Retrieval Prompt and Optimization
-C.1 Aspect Decomposition Prompt
+    End of Annotation Guideline.
+    B.1 Annotator Training
+    The annotation teams each comprised two graduate students.
+To ensure a thorough understanding of bias names and risk levels, all annotators were trained and briefed on their definitions prior to beginning annotation.
+A key part of their preparation involved an in-depth reading of five to ten full-length meta-analyses from Cochrane Train (Main), including their extensive supplementary materials (often over 100 pages), to gain familiarity with relevant terminology and details.
+Although our annotation guidelines allowed for the use of Large Language Models (LLMs) to facilitate understanding of unfamiliar subject matter, annotators infrequently consulted these tools.
+Their preference was to use search engines (with AI overview features disabled).
+On the rare occasions LLMs were employed, their function was strictly confined to assisting with the comprehension of unfamiliar technical details.
+Annotators were monitored and were expressly forbidden from soliciting LLM decisions or opinions regarding annotation choices.
+It is important to note that LLMs were infrequently utilized and only as a supplementary tool for assisting the understanding of unfamiliar subject matters.
+Appendix C Task Descriptions and Visualizations
+We provide visual illustrations of the three tasks.
+C.1 Main Task: Risk-of-Bias Determination
+The model is given the whole biomedical paper, study characteristics (i.e.
+participants, intervention methods, comparators, outcome, etc.), the objective/topic of the systematic review, one bias name and definition, and general guideline that explains how to classify the risk level as high, low or unclear/some concern, and is asked to determine the risk level of bias.
+Note, the definition is based on the specific bias name and risk level.
+Our categorization of different bias names into six main categories does not affect the definition of individual biases.
+See Table 12 for examples of bias categorization.
+Bias Name Bias Category
+random sequence generation (selection bias) selection bias
+bias arising from the randomization process selection bias
+bias due to deviations from intended intervention deviation bias
+bias due to missing outcome data attrition bias
+incomplete outcome data (attrition bias) all outcomes attrition bias
+similarity of other baseline characteristics (selection and performance bias) selection bias / performance bias
+bias in measurement of the outcome detection bias
+blinding of outcome assessment (detection bias) all outcomes detection bias
+blinding for adverse events (performance and detection bias) performance bias / detection bias
+blinding of participants and personnel (performance bias) performance bias
+selective reporting (reporting bias) reporting bias
+bias in selection of the reported result reporting bias
+blinding (performance bias and detection bias) all outcomes performance bias / detection bias
+Table 12: Examples of Bias Categorization
+C.2 Subtask: Support Sentence Retrieval
+When judging the risk of bias for a specific study, the review authors provide support judgments which justify their risk of bias rating.
+These support judgments are grounded in specific aspects of the study, which the support judgment describes.
+This aims to test a model’s ability to retrieve the correct source for the support judgment from a paper’s text.
+Note the model is instructed to retrieve no more than the optimal number of sentences.
+The optimal number can be calculated from Aspect Sentence Mapping and is already included in our dataset.
+See Figure 5 for an illustration.
+C.3 Subtask: Support Judgment Selection
+Figure 6 shows an example of a multiple-choice question with one correct answer and three synthetically generated answers.
+Both options C and D refer to the same information from the paper, in this case the operational challenges affecting the trial’s methodology, but describe different reasoning given this information.
+This demonstrates that retrieving the correct information from the paper is not sufficient for solving this task.
+Figure 5: Subtask: Support Sentence Retrieval.
+In this task, the goal is to retrieve no more than the optimal number of sentences from the biomedical report which support a risk of bias judgment.
+Performance is evaluated against a support judgment, which has been split into aspects.
+Figure 6: Subtask: Support Judgment Selection.
+In this task, the model is shown a paper, and asked which of 7 support judgments provides the best explanation of the paper’s risk of bias.
+Only four options are shown here for illustration purposes.
+Both options C and D refer to the same information from the paper, in this case the operational challenges affecting the trial’s methodology, but describe different reasoning given this information.
+This demonstrates that retrieving the correct information from the paper is not sufficient for solving this task.
+Appendix D Support Sentence Retrieval Prompt, Optimization and Example
+D.1 Aspect Decomposition Prompt
 See Figure 7 for the aspect decomposition prompt.
+D.2 Aspect Filtering Prompt
+See Figure 8 for prompt to filter non-specific commentaries.
 Figure 7: Prompt for decomposing support judgment into aspects.
-C.2 Aspect Filtering Prompt
-See Figure 7 for prompt to filter negations.
-See Figure 8 for prompt to filter non-specific commentaries.
-Figure 7: Prompt for extracting negations.
-Figure 8: Prompt for filtering non-specific commentaries.
-C.3 Aspect-to-Sentence Mapping Prompt and Optimization
-To enhance the agreement between human annotators and the GPT4-0125-turbo model, we developed and tested various prompt engineering strategies on a development set consisting of 30 unique (aspect, paper) pairs.
+D.3 Aspect-to-Sentence Mapping Prompt and Optimization
+See Figure 9 for the prompt optimized on the development set.
+To enhance the agreement between human annotators and the GPT-4-0125-preview model, we developed and tested various prompt engineering strategies on a development set consisting of 30 unique (aspect, paper) pairs.
 It is important to note that no paper or aspect from the development set overlapped with the 50 tasks in the hypothesis testing set.
-GPT4-0125-preview was chosen due to its robust instruction-following capabilities and our familiarity with its performance across different settings.
+GPT-4-0125-preview was chosen due to its robust instruction-following capabilities and our familiarity with its performance across different settings.
 Our experimentation revealed that using in-context learning examples did not improve agreement rates and tended to cause overfitting.
 Instead, we found that embedding the same instruction at both the beginning and the end of the prompt effectively helped the model maintain focus on the task of identifying text elements that significantly cover the details specified in an aspect.
-We also implemented chain-of-thought reasoning, prompting GPT4 to articulate its thought process following a specific keyword.
+We also implemented chain-of-thought reasoning, prompting GPT-4 to articulate its thought process following a specific keyword.
 This approach not only enhanced the quality of the model’s reasoning but also stabilized its performance and reduced common-sense errors.
 To address the issue of entity forgetting in long-context tasks (a typical paper might contain around 5000 tokens), we employed a sliding window technique.
-Each window, containing 10 text elements with a 5-element overlap, allowed GPT4 to process and evaluate each text element within a manageable context size.
+Each window, containing 10 text elements with a 5-element overlap, allowed GPT-4 to process and evaluate each text element within a manageable context size.
 The hyperparameters, window length and overlap size, were optimized using the development set.
@@
 Figure 9 illustrates the final prompt template used to match text elements (such as sentences, tables, figure captions, etc.) with the aspects.
-Each template inputs one aspect, the 10 text elements within a sliding window, and the contextual background (specifically, the summary of the IARC paper from which the aspect was derived).
+Each template inputs one aspect, the 10 text elements within a sliding window, and the contextual background.
 This methodology not only ensures high fidelity in aspect-text alignment but also leverages the model’s capabilities to provide consistent and accurate annotations across extensive text bodies.
-Figure 9: Prompt for generating synthetic options.
-C.4 A Motivating Example
-Here we provide an example of how we build the bias retrieval task.
+D.4 A Motivating Example
+Here we provide an example of how we build the support sentence retrieval task.
 Bias
@@
 Using prompt 8, Aspect 4 is filtered since it is a commentary of the reviewer.
-Mapping Aspects to Sentences in the paper
-Utilizing the procedure described in section C.3, we map the remaining 6 aspects to all sentences in the paper Werneck et al. [2014].
-We only show sentences from the paper that are matched with at least one aspect.
+Mapping Aspects to Sentences in Report
+Utilizing the procedure described in section D.3, we map the remaining 6 aspects to all sentences in the report Werneck et al. (2014).
+We only show sentences from the report that are matched with at least one aspect.
   • Sentence 4: The main outcome is the incidence of infection assessed by the conversion of the Montenegro skin test (MST) after 18 months of follow-up in residents aged $\geq$1 year with no previous history of visceral leishmaniasis (VL).
@@
 (Mapped with aspect 1, 3, 5, 6)
-Bias Recall @ Optimal
-One of our evaluation metric, Bias Recall @ Optimal, measures the the optimal number of sentences to cover all aspects.
+Aspect Recall Ratio @ Optimal
+One of our evaluation metric, Aspect Recall Ratio @ Optimal, measures the the optimal number of sentences to cover all aspects.
 In this example, we only need two sentences, sentence 66 and sentence 73, to cover all aspects.
-Appendix D Task 3 Synthetic Option Generation Procedure and Prompt
+Figure 8: Prompt for filtering non-specific commentaries.
+Figure 9: Prompt for generating synthetic options.
+Appendix E Task-specific Fine-tuning
+We fine-tune Llama-3-8B on the three train sets of RoBBR’s three tasks using LoRA Hu et al.
+(2021), on 1 node of 8 NVIDIA H-100.
+Our fine-tuning hyperparameters are standard: LoRA rank =8=8, LoRA alpha =16=16, a batch size =16=16, with the AdamW optimizer (weight decay =0.01=0.01), and a cosine learning rate scheduler with 10% warmup.
+Llama-3-8B is fine-tuned for 1 epoch using cross entropy loss.
+The model is asked to predict the label (i.e., risk-level for main task, sentence indices for SSR, correct support judgment for SJS).
+We compared different fine-tuning techniques, and found when we train model to predict both the label and a paragraph of reasoning text, the optimal training loss is achieved on the train sets, since it increases the amount of training tokens directly go through cross entropy loss Hsieh et al.
+(2023).
+We use support judgment as proxy for reasoning text.
+Appendix F Analysis: Two Types of Long-Context Problems
+Figure 10: Local-connection problems are easier for LLMs.
+An active research direction in long-context modeling is to categorize different types of long-context problems Li et al.
+(2024).
+Using our annotation pipeline’s Aspect-to-Sentence mapping, we can naturally categorize RoBBR’s datapoints into two types by the optimal number of support sentences used to cover all aspects.
+Local-Connection Problem:
+If one support sentence can cover all aspects that form the support judgment, it means this support sentence is all a model needs to reach the same support judgment.
+While the main task for this datapoint is still a long-context problem because RoBBR’s average document length exceeds 8k tokens, it can be solved by locating one specific sentence from the report.
+There are 213 local-connection datapoints in RoBBR.
+Global-Connection Problem:
+If the support judgment cannot be traced back to one single sentence, some degree of global understanding of the report is required. There are 693 global-connection datapoints in RoBBR.
+In Figure 10, we calculate the average LLM accuracy per type, and found evidence that suggests local-connection problems are easier for both GPT-4o and Sonnet-3.5.
+The intuition is that a problem becomes harder when a more global understanding of the entire long context is required.
+Appendix G Error Analysis
+To better illustrate that GPT-4o and Sonnet-3.5 possess different reasoning strengths, we compare their reasoning traces when both are given the same ground-truth retrieval results.
+Bias: Allocation concealment (selection bias)
+Bias Definition and Selection Criteria for Risk Level:
+Selection bias (biased allocation to interventions) due to inadequate concealment of allocations prior to assignment.
+low risk: Participants and investigators enrolling participants could not foresee assignment because one of the following, or an equivalent method, was used to conceal allocation: Central allocation (including telephone, web-based and pharmacy-controlled randomization); Sequentially numbered drug containers of identical appearance; Sequentially numbered, opaque, sealed envelopes.
+high risk: Participants or investigators enrolling participants could possibly foresee assignments and thus introduce selection bias, such as allocation based on: Using an open random allocation schedule (e.g.
+a list of random numbers); Assignment envelopes were used without appropriate safeguards (e.g.
+if envelopes were unsealed or non-opaque or not sequentially numbered); Alternation or rotation; Date of birth; Case record number; Any other explicitly unconcealed procedure.
+unclear risk: Insufficient information to permit judgement of ’Low risk’ or ’High risk’.
+This is usually the case if the method of concealment is not described or not described in sufficient detail to allow a definite judgement – for example if the use of assignment envelopes is described, but it remains unclear whether envelopes were sequentially numbered, opaque and sealed.
+Retrieved Sentence:
+Patients were randomly assigned to the control or experimental group following a simple randomization procedure (computerized random numbers) achieved using opaque envelopes.
+Physical therapy was started under consultation with the rehabilitation medicine department after wound healing, and the allocation information to each group was not provided to reduce bias.
+Sonnet-3.5:
+REASON: The study describes using a computerized random number generator for randomization, which is a good start for reducing selection bias.
+They also mention using opaque envelopes, which aligns with one of the low-risk criteria (sequentially numbered, opaque, sealed envelopes).
+However, the description doesn’t explicitly state that the envelopes were sequentially numbered or sealed.
+Additionally, while they mention that allocation information was not provided to reduce bias, it’s not entirely clear if this refers to blinding after assignment or concealment before assignment.
+The use of opaque envelopes suggests an attempt at concealment, but without the full details of how these envelopes were used (sequentially numbered, sealed), we can’t be certain that the concealment was adequate to prevent foresight of assignments.
+The computerized randomization is a positive factor, but the implementation of the envelope system is crucial for proper concealment.
+Given these factors, there is some evidence of attempts to conceal allocation, but not enough information to confidently classify it as low risk.
+At the same time, there’s no clear indication of high risk practices.
+DECISION: unclear
+GPT-4o:
+REASON: The study describes the use of opaque envelopes for allocation concealment, which is a method that can prevent participants and investigators from foreseeing assignments, provided that the envelopes are sequentially numbered, opaque, and sealed.
+The highlighted sentences indicate that opaque envelopes were used, but do not specify whether they were sequentially numbered and sealed.
+However, the mention of using opaque envelopes suggests an attempt to conceal allocation, and the additional information that allocation information was not provided to reduce bias further supports the intention to prevent foreseeing assignments.
+Despite the lack of explicit detail about sequential numbering and sealing, the description leans towards a method that is generally considered adequate for allocation concealment.
+DECISION: low
+The provided example demonstrates that Sonnet-3.5 is better in reasoning compared to GPT-4o.
+Sonnet-3.5 correctly identifies critical information gaps in the allocation concealment process—while opaque envelopes were used, there’s no confirmation they were sealed and sequentially numbered, which are important for allocation concealment according to the decision criteria.
+Sonnet-3.5 also recognizes that the mention of withheld allocation information could indicate post-assignment blinding rather than proper pre-assignment concealment.
+In contrast, GPT-4o fails to identify these crucial gaps and makes unwarranted assumptions about concealment adequacy without supporting evidence for key criteria.
+This example clearly illustrates Sonnet-3.5’s better reasoning capabilities.
+Appendix H SJS Synthetic Option Generation Procedure and Prompt
 We first generate three detailed synthetic options that are specifically crafted to imitate the support judgments from other papers concerning the same type of bias, ensuring that they are tailored to be relevant to the specific paper in question while maintaining the foundational reasoning.
 See Figure 11 for the prompt template.
-Following this, we condense these detailed options into shorter versions that preserve the original meaning, using prompt 11.
+Following this, we condense these detailed options into shorter versions that preserve the original meaning, using prompt 12.
 To prevent heuristic algorithms from solving the task easily, we randomly select either the long or short version of each synthetic option to include in the multiple choice questions.
-For selecting the three options derived from other papers’ support judgments within the same bias category, we use prompt 12.
-Finally, once the six incorrect options are constructed, we conduct a manual review with the help of prompt 13 on all data points.
+For selecting the three options derived from other papers’ support judgments within the same bias category, we use prompt 13.
+Finally, once the six incorrect options are constructed, we conduct a manual review with the help of prompt 14 on all data points.
 This review ensures that the options are incorrect and not false negatives.
@@
 Figure 14: Prompt for filtering false negatives.
-Appendix E Experiment Details
-Due to budget limit, we only evaluate the LLMs on a subset of our test set. We provide this subset in our GitHub.
-E.1 Prompt Optimization
+Appendix I Experiment Details
+I.1 Prompt Optimization
 All evaluation prompts are optimized using disjoint development set for each task.
@@
 Consequently, evaluation prompts do not incorporate in-context learning.
-It is important to note that prompt optimization is not tailored to any specific large language model (LLM); instead, the aggregated performance across all models is used to guide the optimization of evaluation prompts.
-E.2 Prompts for Evaluation
-E.2.1 Prompt to Evaluate Study Inclusion/Exclusion (Task 1)
-See Figure 14 for prompt to evaluate Study Inclusion/Exclusion (Task 1).
-Figure 14: Prompt for Study Inclusion/Exclusion (Task1) evaluation
-E.2.2 Prompt to Evaluate Bias Retrieval (Task 2)
-See Figure 15 for prompt to evaluate Bias Retrieval (Task 2). We use an additional multi-turn prompt 16 when the model outputs more than required.
-Figure 15: Prompt for Bias Retrieval (Task 2) evaluation
-Figure 16: Prompt for Bias Retrieval (Task 2) evaluation when the model output more than required
-E.2.3 Prompt to Evaluate Support Judgment Selection (Task 3)
-See Figure 17 for prompt to evaluate Support Judgment Selection (Task 3).
-Figure 17: Prompt for Support Judgment Selection (Task 3) evaluation
-E.2.4 Prompt to Evaluate Risk Level Determination (Task 4)
-See Figure 18 for prompt to evaluate Risk Level Determination (Task 4).
-Figure 18: Prompt for Risk Level Determination (Task 4) evaluation
-E.3 Results with Standard Error
-We provide bootstrapped standard errors for the experimental results 11,12,13,14.
-Table 11: Results for Task 1: Study Inclusion/Exclusion Standard Error Included
-F1 Accuracy Weighted Accuracy
-Claude3-Opus 75.5 $\pm$ 2.6 77.1 $\pm$ 2.5 74.9 $\pm$ 2.5
-Gemini-1.5-Pro 72.2 $\pm$ 2.6 72.7 $\pm$ 2.6 73.0 $\pm$ 2.7
-GPT-4o 75.5 $\pm$ 2.6 75.6 $\pm$ 2.5 77.6 $\pm$ 2.3
-Table 12: Results for Task 2: Sentence Retrieval Standard Error Included
+It is important to note that prompt optimization is not tailored to any specific large language model; instead, the aggregated performance across all models is used to guide the optimization of evaluation prompts.
+I.2 Prompts for Evaluation
+I.2.1 Prompt to Evaluate Support Sentence Retrieval (SSR)
+See Figure 15 for prompt to evaluate Support Sentence Retrieval (SSR). We use an additional multi-turn prompt 16 when the model outputs more than required.
+Figure 15: Prompt for Support Sentence Retrieval (SSR) evaluation.
+Figure 16: Prompt for Support Sentence Retrieval (SSR) evaluation when the model output more than required.
+I.2.2 Prompt to Evaluate Support Judgment Selection (SJS)
+See Figure 17 for prompt to evaluate Support Judgment Selection (SJS).
+Figure 17: Prompt for Support Judgment Selection (SJS) evaluation.
+I.2.3 Prompt to Evaluate Risk-of-Bias Determination (Main Task)
+See Figure 18 for prompt to evaluate Risk-of-Bias Determination (Main Task).
+Figure 18: Prompt for Risk-of-Bias Determination (Main Task) evaluation.
+I.3 Prompt for Risk-of-Bias Determination (Main Task) Given Retrieval Results
+See Figure 19 for prompt for Risk-of-Bias Determination (Main Task) given retrieval results.
+Figure 19: Prompt for Risk-of-Bias Determination (Main Task) Given Retrieval Results.
+I.4 Prompt for OpenAI o4-mini to identify support sentences
+See Figure 20 for prompt for OpenAI o4-mini to identify support sentences
+Figure 20: Prompt for OpenAI o4-mini to identify support sentences.
+I.5 Results with Standard Error
+We provide bootstrapped standard errors for the experimental results in Tables 13,14,15.
 Bias Type
-Model Full Selection Attrition Performance Detection Reporting Analysis Special Other
-n = 337 n = 117 n = 57 n= 50 n = 50 n = 35 n = 13 n = 7 n = 23
-Bias Recall@Optimal
-Claude3-Opus 39.0 $\pm$ 2.4 55.4 $\pm$ 4.0 22.0 $\pm$ 4.7 57.3 $\pm$ 6.39 46.4 $\pm$ 6.5 11.0 $\pm$ 4.9 47.9 $\pm$ 12.3 13.9 $\pm$ 8.3 4.4 $\pm$ 2.9
-Gemini-1.5-Pro 41.1 $\pm$ 2.5 57.2 $\pm$ 4.0 28.7 $\pm$ 5.2 49.1 $\pm$ 6.4 48.8 $\pm$ 6.3 23.3 $\pm$ 6.4 54.2 $\pm$ 11.4 13.9 $\pm$ 8.3 3.3 $\pm$ 2.3
-GPT-4o 44.8 $\pm$ 2.5 60.3 $\pm$ 4.0 28.6 $\pm$ 5.0 64.6 $\pm$ 5.5 48.7 $\pm$ 6.4 32.9 $\pm$ 7.3 28.7 $\pm$ 9.7 28.0 $\pm$ 14.1 2.3 $\pm$ 2.2
-Bias Recall@3
-Claude3-Opus 52.3 $\pm$ 2.5 67.9 $\pm$ 3.8 40.2 $\pm$ 5.8 63.1 $\pm$ 5.9 55.9 $\pm$ 6.6 25.5 $\pm$ 7.1 55.9 $\pm$ 12.3 64.6 $\pm$ 16.9 19.7 $\pm$ 7.8
-Gemini-1.5-Pro 56.1 $\pm$ 2.4 69.7 $\pm$ 3.8 44.1 $\pm$ 5.7 73.7 $\pm$ 5.3 65.3 $\pm$ 6.1 35.6 $\pm$ 7.6 56.3 $\pm$ 12.8 35.5 $\pm$ 16.1 6.7 $\pm$ 4.8
-GPT-4o 61.8 $\pm$ 2.4 72.2 $\pm$ 3.8 47.9 $\pm$ 5.8 82.0 $\pm$ 4.5 78.5 $\pm$ 5.4 43.2 $\pm$ 8.2 67.9 $\pm$ 10.3 50.4 $\pm$ 17.1 10.3 $\pm$ 5.8
-Table 13: Results for Task 3: Support Judgment Selection Standard Error Included
+Model Avg Selection Attrition Performance Detection Reporting
+n = 157 n = 46 n= 54 n = 61 n = 14
+OpenAI-v3 22.72 40.13 $\pm$ 3.56 6.3 $\pm$ 2.72 26.67 $\pm$ 5.29 27.38 $\pm$ 4.81 13.1 $\pm$ 7.56
+GritLM-7B 18.87 35.84 $\pm$ 3.59 6.99 $\pm$ 2.67 15.19 $\pm$ 3.76 26.83 $\pm$ 5.08 9.52 $\pm$ 7.09
+GPT-4o 47.48 60.5 $\pm$ 3.54 42.36 $\pm$ 6.47 41.48 $\pm$ 6.06 50.05 $\pm$ 5.7 43.03 $\pm$ 11.55
+Sonnet-3.5 39.17 55.42 $\pm$ 3.62 30.65 $\pm$ 5.71 37.01 $\pm$ 5.69 37.21 $\pm$ 5.5 35.54 $\pm$ 11.55
+Llama-3.1-70B 45.62 61.17 $\pm$ 3.48 34.21 $\pm$ 5.95 45.19 $\pm$ 5.96 50.11 $\pm$ 5.68 37.41 $\pm$ 11.49
+Llama-3-8B 22.66 49.44 $\pm$ 3.68 14.49 $\pm$ 4.86 22.13 $\pm$ 4.95 20.77 $\pm$ 4.7 6.46 $\pm$ 4.4
+Llama-3-8B Fine-Tuned 40.8 69.01 $\pm$ 3.41 24.78 $\pm$ 5.85 47.28 $\pm$ 6.26 48.63 $\pm$ 5.92 14.29 $\pm$ 9.35
+Table 13: Results for Sentence Retrieval Standard Error Included.
 Bias Type
-Model Full Selection Attrition Performance Detection Reporting Analysis Special
-n = 310 n = 86 n = 60 n = 52 n = 56 n = 50 n = 8 n = 6
-Accuracy
-Claude3-Opus 51.2 $\pm$ 2.8 62.6 $\pm$ 5.2 56.7 $\pm$ 6.4 48.0 $\pm$ 6.9 46.5 $\pm$ 6.9 44.2 $\pm$ 7.0 50.0 $\pm$ 17.9 0.0 $\pm$ 0.0
-Gemini-1.5-Pro 49.3 $\pm$ 3.0 57.1 $\pm$ 5.3 59.9 $\pm$ 6.5 46.4 $\pm$ 7.0 42.8 $\pm$ 6.8 40.2 $\pm$ 7.0 49.3 $\pm$ 17.7 16.6 $\pm$ 15.5
-GPT-4o 45.8 $\pm$ 2.9 52.3 $\pm$ 5.4 63.1 $\pm$ 6.2 42.5 $\pm$ 6.7 37.7 $\pm$ 6.5 28.0 $\pm$ 6.2 74.5 $\pm$ 15.7 0.0 $\pm$ 0.0
-Table 14: Results for Task 4: Risk Level Determination Standard Error Included
+Model Full Selection Attrition Performance Detection Reporting
+n = 310 n = 86 n = 60 n = 52 n = 56 n = 50
+GPT-4o 47.17 58.46 $\pm$ 4.37 60.2 $\pm$ 5.04 48.28 $\pm$ 5.24 42.68 $\pm$ 5.29 26.25 $\pm$ 4.7
+Sonnet-3.5 59.92 73.08 $\pm$ 3.89 73.47 $\pm$ 4.28 50.57 $\pm$ 5.42 51.22 $\pm$ 5.44 51.25 $\pm$ 5.26
+Llama-3.1-70B 53.16 66.15 $\pm$ 3.89 62.24 $\pm$ 4.98 44.83 $\pm$ 5.19 46.34 $\pm$ 5.64 46.25 $\pm$ 5.7
+Llama-3-8B 26.54 26.92 $\pm$ 3.87 34.69 $\pm$ 4.77 24.14 $\pm$ 4.64 21.95 $\pm$ 4.69 25.0 $\pm$ 4.91
+Llama-3-8B Fine-Tuned 29.6 40.77 $\pm$ 4.36 28.57 $\pm$ 4.37 24.14 $\pm$ 4.42 19.51 $\pm$ 4.52 35.0 $\pm$ 5.38
+Table 14: Results for Support Judgment Selection Standard Error Included.
 Bias Type
-Model Full Selection Attrition Performance Detection Reporting Analysis Special Other
-n = 337 n = 93 n = 50 n = 50 n = 50 n = 50 n = 18 n = 26 n = 50
-F1
-Claude3-Opus 47.7 $\pm$ 2.9 54.3 $\pm$ 6.6 34.1 $\pm$ 7.0 46.7 $\pm$ 6.8 37.4 $\pm$ 7.6 23.0 $\pm$ 4.7 69.9 $\pm$ 18.8 27.9 $\pm$ 6.7 27.9 $\pm$ 1.8
-Gemini-1.5-Pro 43.1 $\pm$ 2.6 53.9 $\pm$ 5.8 26.0 $\pm$ 3.7 43.7 $\pm$ 7.2 46.8 $\pm$ 7.4 30.4 $\pm$ 5.2 44.6 $\pm$ 7.7 20.1 $\pm$ 3.4 22.8 $\pm$ 6.1
-GPT-4o 47.4 $\pm$ 2.9 54.3 $\pm$ 6.6 30.4 $\pm$ 5.8 46.7 $\pm$ 6.8 37.4 $\pm$ 7.6 22.6 $\pm$ 3.9 69.9 $\pm$ 18.8 32.2 $\pm$ 8.5 27.0 $\pm$ 1.6
-Accuracy
-Claude3-Opus 56.8 $\pm$ 2.6 64.6 $\pm$ 5.0 58.1 $\pm$ 7.1 48.2 $\pm$ 7.1 41.7 $\pm$ 7.1 40.0 $\pm$ 7.3 88.9 $\pm$ 7.3 53.8 $\pm$ 9.7 72.2 $\pm$ 6.4
-Gemini-1.5-Pro 49.3 $\pm$ 2.7 62.5 $\pm$ 5.1 42.4 $\pm$ 7.2 46.5 $\pm$ 6.8 48.1 $\pm$ 7.1 56.0 $\pm$ 7.2 77.5 $\pm$ 9.5 42.4 $\pm$ 9.9 23.7 $\pm$ 6.2
-GPT-4o 57.0 $\pm$ 2.6 64.6 $\pm$ 5.0 58.1 $\pm$ 7.2 48.2 $\pm$ 7.1 41.7 $\pm$ 7.2 43.6 $\pm$ 7.2 88.9 $\pm$ 7.3 57.7 $\pm$ 9.7 68.3 $\pm$ 6.5
-Weighted Accuracy
-Claude3-Opus 49.5 $\pm$ 3.0 55.3 $\pm$ 5.1 36.2 $\pm$ 9.0 48.0 $\pm$ 7.0 46.5 $\pm$ 5.9 30.4 $\pm$ 10.1 69.3 $\pm$ 19.6 28.2 $\pm$ 4.9 32.4 $\pm$ 0.9
-Gemini-1.5-Pro 47.9 $\pm$ 3.2 56.3 $\pm$ 5.6 34.0 $\pm$ 10.1 44.4 $\pm$ 7.0 54.2 $\pm$ 8.2 36.4 $\pm$ 10.1 43.6 $\pm$ 4.2 22.2 $\pm$ 4.9 32.5 $\pm$ 10.1
-GPT-4o 49.3 $\pm$ 3.1 55.3 $\pm$ 5.1 29.3 $\pm$ 3.9 48.0 $\pm$ 7.0 46.5 $\pm$ 5.9 24.5 $\pm$ 8.4 69.3 $\pm$ 19.6 60.4 $\pm$ 24.4 30.6 $\pm$ 1.5
-Appendix F Full Example
-We provide the full version of the example presented in the main paper. The example is shown in its entirety here, whereas in the main paper it was shortened to save space.
-F.1 Search Protocol Gonzalez et al. [2015]
-Types of Studies Randomized controlled trials (RCTs)
-Types of Participants People living in leishmaniasis endemic regions.
-Types of Interventions Any intervention that aims to reduce leishmaniasis incidence through vector or reservoir control.
-Types of Outcome Measures People developing CL or VL infections.
-Estimates of the vector density measured by an appropriate technique (adult sandfly density estimated by counts of vectors either landing on exposed body parts of humans acting as baits or collected resting inside buildings, for example, on walls).
-Number of participants with positive immunological or biochemical tests that detect contact with the parasite (for example, leishmanin skin test conversion rates or lymphocyte proliferation rates, or both).
-Adverse effects on people.
-Adherence to control measures; for example, the extent to which specified intervention components were delivered as prescribed.
-Measures of environmental impact (assessment of the possible impact - positive or negative - that the interventions may have on the natural environment) or sustainability (assessment of the ability to change biological and human processes, functions, biodiversity and productivity), or both.
-F.2 Objective of the Meta-Analysis Gonzalez et al. [2015]
-To assess the effects of vector and reservoir control interventions for cutaneous and for visceral leishmaniasis.
-F.3 Full Paper
-See Werneck et al. for full paper content.
-F.4 Bias Name and Definition Higgins and Green [2011]
-Bias Name
-selective reporting (reporting bias)
-Bias Definition
-Reporting bias due to selective outcome reporting.
-F.5 PICO
-Methods Trial design: Cluster-RCT.
-Unit of randomization: Geographic area.
-Number of clusters: 40 geographic areas.
-Entomological data collection: Not done.
-Clinical data collection: Conversion of the Montenegro skin test (MST) at 18 months of follow-up.
-Length of follow-up: 18 months.
-Analysis: Analysed at cluster level.
-Participants Ten localities in 7 neighbourhoods of the city of Teresina (Brazil) were divided into blocks, each containing an average of 60 residences.
-For each locality, 4 blocks were selected to minimize the risk of cross-contamination of interventions.
-Eligible participants were residents of selected blocks aged 1 year or above with no history of VL.
-The 40 geographic areas (blocks) randomly allocated to the 4 types of interventions (697 subjects MST-).
-Endemic disease: VL caused by L.
-chagasi (L.
-infantum).
-Interventions 1.
-Spraying households and residential annexes with insecticide.
-2.
-Elimination of infected dogs.
-3.
-Combination of spraying and eliminating infected dogs.
-4.
-No intervention.
-Description of spraying: performed according to the routine of the VL Control Program of the Zoonosis Control Center of the Teresina City Health Department.
-Interventions were delivered in the selected blocks every 6 months, for three times, beginning just after each household visit.
-The elimination of infected dogs was decided if indirect immunofluorescence test was more or equalled 1:40.
-Outcomes Cases of infection by L. infantum at 18 months determined by conversion of the MST (MST- at the beginning) or diagnosis of active VL.
-Notes Country: Brazil (Teresina, Itararé quarter).
-Trial dates: January 2004 to December 2006.
-Trial sponsor: Funded by Health Surveillance Unit from the Brazilian Ministry of Health.
-One author was partially funded by the Brazilian Research Council (CNPq 306267/2010-1 and 202088/2012-0).
-The founders had no role in trial design, data collection and analysis, decision to publish, or preparation of the manuscript.
-The authors have declared that no competing interests exist.
-Sample size: Calculated.
-Compliance assessment: Not reported.
-[…]
-[…]
+Model Avg Selection Attrition Performance Detection Reporting Deviation
+n = 933 n = 629 n = 309 n = 645 n = 594 n= 331
+GPT-4o 42.07 50.52 $\pm$ 1.87 36.48 $\pm$ 2.14 54.67 $\pm$ 2.8 43.66 $\pm$ 2.02 34.43 $\pm$ 1.87 32.65 $\pm$ 2.77
+Sonnet-3.5 41.93 52.83 $\pm$ 1.86 37.07 $\pm$ 2.04 50.56 $\pm$ 2.75 43.29 $\pm$ 2.09 33.58 $\pm$ 1.41 34.23 $\pm$ 1.96
+Llama-3.1-70B 38.81 48.1 $\pm$ 1.85 31.7 $\pm$ 1.94 48.04 $\pm$ 2.67 44.36 $\pm$ 2.0 31.58 $\pm$ 1.33 29.04 $\pm$ 2.37
+Llama-3-8B 30.05 36.37 $\pm$ 1.28 32.44 $\pm$ 1.62 39.13 $\pm$ 2.59 37.21 $\pm$ 1.92 19.76 $\pm$ 1.96 15.38 $\pm$ 1.64
+Llama-3-8B Fine-Tuned 36.33 49.51 $\pm$ 1.91 33.34 $\pm$ 1.89 42.44 $\pm$ 2.69 40.1 $\pm$ 2.01 26.95 $\pm$ 0.44 25.67 $\pm$ 1.06
+Table 15: Results for Risk Level Determination Standard Error Included.